A Complete Bibliography of Publications in Reviews of Modern Physics (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.05

Title word cross-reference

2 [Pos06]. 5f [MvdL09]. ^{+} [BUS08]. ^{2} [DW08]. ^{3} [DP02, Leg04]. ^{4} [GKMDA08a, GKMDA08b, WHG01]. _{2} [MM03]. _{3} [CFCT09, JT02]. _{4} [MM03]. _{60} [CFCT09]. _{c} [BT05, HM02, KW09]. _{L} [BUS08]. β [Zde02]. d [Ste01, Ste06]. ϵ′/ϵ [BFE00]. f [Pfl09, SCA^{+}09, Ste01, Ste06]. H [SKC02]. M [Tay01]. μ [SBK00]. → π^{+}νν [BUS08]. → π^{0}νν [BUS08]. SO(5) [DHZ04]. Z [RDR06, RDM00, Lan09].

-electron [Pfl09, SCA^{+}09, Ste01, Ste06].

70 [Gre00]. 72 [Zan01, ’t 02]. 73 [Leg03]. 76 [BFK^{+}04b, BFK^{+}04a]. 78 [GHL^{+}09, JTLJ06a]. 79 [HKP^{+}07a, KMN^{+}07b, PNHRD09].

80 [GKMDA08b]. 81 [EHM14, Mei09b, Nam10].

Abelian [NSS^{+}08]. absorption [RA00]. ac [DS00]. accelerators [ESL09]. accumulators [Wei03]. achieve [RGC^{+}08]. acoustic [PLT02]. Acoustical [May01]. actinide [MvdL09, SCA^{+}09]. Addendum [Ste06]. adiabatic [KTS07]. Advances [Gie03]. Aligning [SS03]. alkali [Leg01, Leg03]. alloy [vdWC02]. America [RGC^{+}08]. American [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. amorphous [PLT02]. analog [DC08]. analogs [May01]. Anderson [EM08]. Andreev [Bee08, Deu05]. Angle [DHS03]. Angle-resolved [DHS03]. anisotropies [Smo07]. annealing [DC08]. Announcement [Ano07]. antiferromagnetism [DHZ04]. antimatter [DK04]. anyons [NSS^{+}08]. Application [Gre98, Gre00]. applications [Alf01, CHM08, GHL^{+}06, GHL^{+}09, KK08, OW07, RDB^{+}09, ZFS04]. approach [BM03, DKI^{+}03]. approaches [NM02, ORR02, RA00, ZD08]. APS [Spr08a]. argon [GG08]. arrays [SHS06, dA07]. Artificial [HM09]. Aspects [Cre01, GMD02]. astronomy [Gia03, Sah02]. astrophysical [ZMD04]. astrophysics [KP07, Kos03, RDR06]. asymmetry [DK04]. Asymptotic [Wil05, Gro05]. atmospheric [KT01]. atom [BMH08, Ket02]. Atomic [KP07, Gie03, GPS08, HFS03, JTLJ06a, JTLJ06b, Luk03, Wie09]. atoms [BGK^{+}02, CSP09, FZ07, Ket02, RBH01, TRR00]. atrix [Tay01]. attenuation [PLT02]. Attosecond [KI09]. attribution [Pol05]. authors [Spr08a]. avalanches [AJ04]. average [NM02]. axions [BCK^{+}03].

B [Hur03]. background [Smo07]. band [KIC^{+}08, Kro01]. Bang [Mat07]. based [BGV02, ESL09, RWP05]. beams [Wei03]. beginning [Gin04]. behave [Ket02]. behavior [AKS01, AT06, Ste01, Ste06]. Bénard [AGL09]. beta [AEE08, SBNC06]. between [KMM09, TRR00]. beyond [Grü08, KO01, Mat07]. Big [Mat07]. biological [GNS02, KLLW07]. Birth [Kos03]. body [AFOV08, BMH08, BDZ08, DPS04, KIC^{+}08, ORR02]. Bogoliubov [OKSD05]. Boltzmann [SKC02]. Boost [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. Boost-Phase [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. Bose [And04, CW02, Fet09, Ket02, Leg01, Leg03, MO06, OKSD05, Pos06]. boson [TKKTBW01]. bosons [Lan09]. boundaries [HM02]. boundary [VG07]. bounded [PLB08]. branched [JL09]. breaking [Fra01, Nam09, Nam10, SS00]. brown [BHLL01]. Brownian [Gre07, HM09]. bubble [BHL02, Mor07]. Bulk [OKSD05, Str09]. bundles [CCG^{+}01]. bursts [Pir04].

C [CFCT09]. calculations [KSH^{+}06]. Calorimetry [FG03]. can [RGC^{+}08]. carbon [LML02]. Carlo [FMNR01]. case [Nam09, Nam10]. Casimir [KMM09]. cavity [BCK^{+}03, RBH01]. center [CB09]. century [Dav03, Gin04]. challenge [vTS00]. chaos [PW07, WM09]. Chaotic [Sha08]. Charge [BKM02, GNS02]. Charge-transfer [BKM02]. chemical [GNS02]. chemistry [BM07]. Chern [Mar05]. Chiral [Cha03, Cre01]. classes [Ódo04]. Classical [GMD02, BKV05, Zur03]. cloning [SIGA05]. closed [OW07]. cluster [BM07, MJPH05]. clusters [Mor07, Voi05]. CODATA [MT00, MT05, MTN08]. coherent [FIM05]. Coherently [KTS07]. cold [KGJ06]. collective [AT06, MS06]. colliders [RZ00, TKKTBW01]. collisionally [RWP05]. collisions [BMH08]. Colloquium [AJ04, AP06, Bad06, BM03, Bat07, Bee08, BBF^{+}08, BMW09, CFCT09, CY03, DC08, DPS04, Dyr06, ECS04, FH03, Fie07, FT03, GLP^{+}02, GPM06, Gon03, GNS02, GCH03, JL09, KTS07, KW09, LW05, Luk03, MNV09, Mor07, Ost05, OKSD05, PW07, PNHRD07, Pos06, PLB08, RDB^{+}09, RWP05, Sac03, SHS06, Sha08, SE09, Sid09, SS03, Str09, SKC02, Tho09, Tso09, WSG08, YR09, Zde02, ZMD04, ZD08, dA07, PNHRD09]. Color [ASRS08, Ano07]. combs [CY03]. compactification [DK07]. complete [TRR00]. Complex [MI09, AB02, AK02, DGM08]. compounds [MRMN03, Pfl09]. comprehensive [DKI^{+}03]. computation [DC08, GMD02, NSS^{+}08, VC04]. computing [KMN^{+}07a, KMN^{+}07b]. concept [Alf01]. concepts [AT06, Leg01, Leg03, RDB^{+}09]. condensate [OKSD05]. condensates [Fet09, MO06]. condensation [CW02, Ket02, Leg01, Leg03]. Condensed [CCG^{+}01, CN04, May01]. condensed-matter [May01]. conditions [WHG01]. conductance [ECS04]. conducting [Shi01]. conduction [DS00]. conductivity [PLT02]. configuration [Kas08]. confined [Boo04]. conformations [JL09]. confrontation [’t 00, ’t 02]. Conifolds [GK08]. consequences [DRW06]. consistency [GG08]. consistent [BHR03]. constant [PR03]. constants [MT00, MT05, MTN08, Uza03]. Continuous [LR09, BvL05, KF08]. Continuous-variable [LR09]. Control [LK04, Bec05, VC04, WSG08]. controlled [KTS07]. Controlling [HM09, Str09]. convection [AGL09]. conventional [BVZ06, MZVW01]. copyright [Spr08a]. core [BKM02]. corrals [FH03]. correlated [ABGH09, DH07, Sha08]. Cosmic [Smo07, NW00, Voi05]. cosmological [PR03]. cosmology [BW06]. counting [EHM09b, EHM14]. Coupled [BM07]. Coupled-cluster [BM07]. coupling [Das04, GF04]. CP [BFE00, IN08, Kob09, Mas09]. CP-violating [BFE00]. Critical [BHLD07, DGM08]. Criticality [WHG01]. cross [Nam09, Nam10]. cryptography [GRTZ02]. crystal [BdGDG01, DGC06, GPM06, IS06]. crystalline [Ugg05]. crystallization [Str09]. crystallography [RRG01]. crystals [BAP05, vTS00]. cuprate [DHS03, Deu05, Sac03, TK00]. current [GKI04]. current-phase [GKI04]. curves [OW07]. cycle [BK00]. cyclotron [Chu04].

D [Pos06]. dark [BCK^{+}03, PR03]. dark-matter [BCK^{+}03]. Data [GG08, KP07]. dawn [Gia03]. dawning [Shi01]. days [Leg04]. decay [AEE08, KO01, SBNC06, Zde02]. decays [DHZ06, Hur03]. Decoherence [Sch04, Zur03]. Defects [ABGH09, KF08]. Defense [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. Defining [Hal06]. demon [MNV09]. dense [ASRS08]. density [BdGDG01, Bec00, GR09, KK08, ORR02, Sch05]. density-functional [BdGDG01, Bec00, ORR02]. density-matrix [Sch05]. dependent [KK08]. deposition [WH00]. derivative [Spr08a]. description [Tho09]. design [PGT00]. detect [KBF^{+}03]. detection [ZD08]. determination [LPT03]. development [Fer08]. developments [OS00]. devices [Gre98, Gre00, MSS01]. diagram [BMW09]. dielectric [PNHRD07, PNHRD09]. diffraction [LML02]. Diffusion [FFM^{+}03]. dilemma [Pol05]. dilute [CW02, Pos06]. dimensional [DMW00]. dimensions [AKS01, KM07]. dioxides [SCA^{+}09]. Dipolar [DMW00]. Dirac [Bat07]. disaster [JBWS00]. discharge [LLLG04]. Disclinations [KF08]. discovery [Gro05, HM00, Shi01, Smo07]. dislocations [KF08]. disordered [DS00, Sid09]. Dissipation [KM07]. Dissipation-induced [KM07]. distribution [JMST04, SBPC^{+}09]. DNA [ECS04, Vio00, ZD08]. do [Gin04]. does [Mas09]. dominated [RWP05]. Doping [LNW06]. dots [Alh00, HKP^{+}07a, HKP^{+}07b, RM02, Sha08, vdWDE^{+}03]. Double [AEE08, Alf01, Zde02, vdWDE^{+}03]. drag [PLB08]. driven [ESL09, Hel01]. duality [SSJ04]. dust [SE09]. dust-plasma [SE09]. dwarfs [BHLL01]. Dynamical [RVSA06, SS00, FZ03, KSH^{+}06]. Dynamics [CM09, MO06, AVS09, AGL09, BTW06, BKM02, CFL09, FS00, LBMW03, SRS05, TBBH05]. dynamos [GLP^{+}02].

early [Leg04]. edge [Cha03]. Editorial [Spr08a, Spr08b, Spr08c]. effect [Bat07, CHM08, MS03, vdWC02]. Effective [BPSV05]. Effective-field [BPSV05]. effects [BF05, BGK^{+}02, Buz05, CT05, DMW00, KIC^{+}08, MS06]. einselection [Zur03]. Einstein [Leg03, CW02, Fet09, Ket02, Leg01, MO06, OKSD05, RDB^{+}09]. elastic [Dyr06, Wit07]. electric [Tso09]. electrical [GCH03]. Electrodynamics [BT05]. electromagnetic [PNHRD07, PNHRD09]. Electromagnetically [FIM05]. Electron [GG08, KW09, vdWDE^{+}03, Bat07, BDS08, Chu04, ESL09, HKP^{+}07a, HKP^{+}07b, LML02, NM02, Pfl09, SCA^{+}09, Ste01, Ste06, TRR00, Tho09]. Electron-lattice [KW09]. electron-nucleus [BDS08]. electron-positron [Tho09]. Electronic [CBR07, GPM06, KSH^{+}06, ORR02, RM02, BLVE07, NGP^{+}09]. electronics [Alf01]. electrons [DH07, KS01, Kro01, LW05, Sha08]. Electrophoresis [Vio00]. Electrostatic [ABD^{+}06]. electroweak [SBNC06]. elements [HM00]. emergence [Gro05]. emergent [LW05]. Energetic [BF05]. Energetics [Zan00, Zan01]. energies [DKI^{+}03]. energy [BMH08, GR09, NW00, PR03, RGC^{+}08, SHS06]. engineering [MSS01]. ensembles [Luk03]. Entanglement [AFOV08, HHHH09, RBH01]. entropy [Ved02]. environment [Fer01]. equation [AK02]. era [Shi01]. Erratum [EHM14, GHL^{+}09, Gre00, Leg03, PNHRD09, Zan01, ’t 02]. essay [Bec05]. evaluation [DKI^{+}03, GG08]. evidence [GGN03]. evolution [Voi05, WHW02]. Exact [FZ03]. Exactly [DPS04]. exceptional [Spr08b]. excitations [ORR02, OKSD05]. expanded [CFCT09]. Experiment [KMM09]. Experimental [RDR06, SRS05]. Experiments [AJ04, BGV02, CW02, GLP^{+}02]. explosion [WHW02]. extragalactic [Wid02]. extrasolar [BHLL01].

Factory [BGP^{+}09]. Falicov [FZ03]. fast [RDM00]. Feedback [Bec05]. Femtosecond [CY03, RRG01]. Fermi [GPS08, vLRVW07, Ste01, Ste06]. Fermi-liquid [vLRVW07]. fermions [FS00]. ferroelectric [DRS05]. ferromagnet [BVE05, Buz05]. ferromagnetic [JSM^{+}06, TBBH05]. ferromagnets [KIC^{+}08]. fertilization [Nam09, Nam10]. Feshbach [KGJ06]. few [BK00, HKP^{+}07a, HKP^{+}07b]. few-cycle [BK00]. few-electron [HKP^{+}07a, HKP^{+}07b]. fiber [DGC06]. Field [Tho09, BHR03, BPSV05, DN01, FZ03, KSH^{+}06, MI09, VP07]. fields [BK00, FGV01, Kro01, Lai01, Tso09, Ugg05, Wid02]. Figures [Ano07]. film [DRS05, Shi01]. films [CM09, DMW00]. Final [RZ00]. Final-focus [RZ00]. fine [RA00]. Finite [GKMDA08a, GKMDA08b, DHJ03, KM07]. Finite-size [GKMDA08a, GKMDA08b]. first [CW02]. Flavor [BGP^{+}09, Kob09]. flow [Gro00, Ter00]. flows [VG07]. fluctuating [KBF^{+}03]. fluctuation [EHM09b, EHM14]. fluctuations [EHM09b, EHM14, dJOS03]. Fluid [SQ05, FGV01, RWP05]. Flux [DK07]. focus [RZ00]. focusing [Wit07]. folding [BM03, PGT00]. force [Gie03, KMM09]. forces [EHM09a]. formation [LK04]. forming [Dyr06]. Foundations [BH07]. Four [BMH08]. Four-body [BMH08]. fourth [Hee01]. Fractal [AVS09]. fractional [Cha03, MS03]. fragmentation [TRR00]. frames [BRS07]. free [NM02]. free-electron [NM02]. freedom [Gro05, Wil05]. freezing [PGT00]. frequencies [Hal06]. frequency [CY03]. friction [VP07]. Frontiers [BK00]. fullerides [CFCT09]. function [ORR02]. functional [BdGDG01, Bec00, ORR02]. functionals [KK08]. fundamental [Leg01, Leg03, MT00, MT05, MTN08, Tay01, Uza03]. Fundamentals [SE09, ZFS04]. future [BHLD07, Fer08, Zde02].

galactic [Wid02]. galaxies [HDB^{+}09, JMST04, Voi05]. galaxy [Fer01]. gamma [Pir04]. gamma-ray [Pir04]. gas [And04, BW06, CW02]. gases [BDZ08, CCG^{+}01, GPS08, Leg01, Leg03, Pos06]. Gaudin [DPS04]. Gauge [OS00, JO01, Lan09]. generation [DGC06, Hee01, WSG08]. generic [BKV05]. Geodynamo [RG00]. geometric [GK08]. Geometrical [BM03]. geometry [Kam02]. geophysical [DRW06]. giant [BHLL01, Grü08]. Ginzburg [AK02]. glass [Das04, Dyr06]. glass-forming [Dyr06]. glasses [FFM^{+}03]. global [RGC^{+}08]. Grain [HM02]. Granular [BLVE07, AT06]. graphene [Bee08, NGP^{+}09]. gravitation [Vel00]. Green’s [ORR02]. Green’s-function [ORR02]. ground [TRR00]. Group [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c, BCP08, MRMN03, Sch05, VG07]. group-IV [MRMN03]. guiding [CB09]. guiding-center [CB09]. gyrokinetic [BH07].

Hadron [DW08, TKKTBW01]. hadronic [DHZ06]. Half [KIC^{+}08, Dav03]. half-century [Dav03]. Half-metallic [KIC^{+}08]. Hall [Cha03, MS03]. halos [JRFG04]. Hamiltonian [CB09, MS03]. hard [Kla02]. harmonic [WSG08]. harmonics [TG09]. Heat [AGL09, VP07]. heaviest [HM00]. heavy [BPSV05, Lan09]. helices [KLLW07]. Helioseismology [CD02]. helium [BAP05]. Heteropolymer [PGT00]. heterostructure [Alf01]. heterostructures [Buz05, IS06, TBBH05]. High [MRMN03, TG09, BT05, BMH08, ECS04, FKMA^{+}07, GR09, HM02, KBF^{+}03, KW09, LNW06, NM02, RDR06, Wei03, WSG08]. high-average-power [NM02]. high-conductance [ECS04]. high-energy [BMH08]. high-harmonic [WSG08]. high-intensity [Wei03]. High-order [TG09]. High-pressure [MRMN03]. high-T [BT05, HM02]. high-temperature [FKMA^{+}07, KBF^{+}03, LNW06]. Historical [JO01, OS00]. hole [dA07]. holographic [Bou02]. hundred [Gla06]. hydrodynamic [ES06, SKC02]. hydromagnetic [GLP^{+}02]. hydrophobicity [AP06].

ice [DRW06]. II [Abr04, MZVW01, MRMN03, SBK00]. III [JSM^{+}06, MRMN03]. Illuminating [Bat07]. impact [KW09, Sah02]. impedance [Sid09]. implications [GGN03, NW00]. Impurity [BVZ06, BCP08]. Impurity-induced [BVZ06]. Inclusive [BDS08, Hur03]. incoherent [Fie07]. income [YR09]. Induced [GF04, BVZ06, FIM05, KM07]. Inelastic [DH07, KS01]. infinity [’t 00, ’t 02]. Inflation [BTW06]. influences [BKV05]. Information [GMD02, BRS07, BvL05, MNV09, PT04, Ved02]. Inhomogeneous [CN04]. instabilities [KM07, vLRVW07]. insulator [LNW06]. Intense [BK00]. intensity [Wei03]. interacting [And04, BMW09, Pos06]. interaction [GF04, KW09, LMP03, Ugg05]. interactions [JL09, KLLW07, MS06, SCA^{+}09, SE09, Vel00]. Intercept [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. interdisciplinary [MI09]. interfaces [IS06]. interference [Gre98, Gre00]. interferometry [CSP09, Sah02]. intergalactic [Mei09a, Mei09b]. interpretations [Sch04]. interstellar [Fer01]. invariance [BKV05, JO01]. inversion [GNS02]. ion [BMH08, Sid09]. ion-atom [BMH08]. ionized [HDB^{+}09]. ions [LBMW03]. irradiated [TG09]. Isotope [CT05]. Issues [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. IV [MRMN03].

James [Deu05]. Josephson [DP02, GKI04, MSS01]. Josephson-junction [MSS01]. Journal [Ano07]. junction [MSS01]. junctions [GKI04]. just [Hog00].

Kapitza [Bat07]. kernel [WWAF06]. key [SBPC^{+}09]. Kimball [FZ03]. kinetic [BF05]. Klein [Bee08]. Kuramoto [ABV^{+}05].

Laboratory [GLP^{+}02]. lamps [LLLG04]. Landau [AK02]. large [AGL09]. laser [BK00, ESL09, Ket02, SS03, TG09, WH00]. laser-driven [ESL09]. laser-irradiated [TG09]. lasers [NM02, RDR06]. lattice [Abr04, Cre01, KW09, Ódo04, SKC02, vdWC02]. lattices [MO06]. Laws [Str09, JMST04]. layers [BDV07, VG07]. leaves [Spr08a]. Lecture [Nam10, ’t 02, Abr04, Alf01, CW02, Dav03, Fer08, Gia03, Gin04, Gla06, Gro05, Grü08, Hal06, Hän06, Hee01, Ket02, Kob09, Kos03, Kro01, Leg04, Mac01, Mas09, Mat07, Nam09, Pol05, Shi01, Smo07, Vel00, Wil05, ’t 00]. length [AP06]. Light [dA07, DH07, Gla06]. Linear [KMN^{+}07a, RZ00, KMN^{+}07b]. links [DP02]. liquid [CM09, Fie07, vLRVW07, SRS05, Ste01, Ste06, WHG01, vTS00]. liquids [Cha03, Das04, Dyr06, Tso09]. localization [SHS06]. Long [JTLJ06a, JTLJ06b]. Long-range [JTLJ06a, JTLJ06b]. look [RGC^{+}08]. Low [PLT02, DW08, RWP05]. Low-temperature [PLT02, RWP05]. Luttinger [Cha03, Fie07].

Magnetic [FZ07, DMW00, Gre98, Gre00, Lai01, vLRVW07, MZVW01, Wid02]. magnetically [Boo04, KGJ06]. magnetization [TBBH05]. magneto [BGK^{+}02]. magneto-optical [BGK^{+}02]. Magnetohydrodynamic [ZMD04]. magnetoresistance [Grü08]. Majorana [AEE08]. managed [Gin04]. manganites [SJ01]. Manipulating [RBH01, Luk03]. manipulation [SHS06]. Many [BDZ08, AFOV08, DPS04, Hel01, KIC^{+}08, ORR02]. Many-body [BDZ08, AFOV08, DPS04, KIC^{+}08, ORR02]. many-particle [Hel01]. mapping [Wie09]. maser [Chu04]. mass [AEE08]. masses [GGN03, LPT03]. massive [WHW02]. materials [ABD^{+}06, Hee01, Kam02, KMM09]. matrices [PW07, WM09]. matrix [Sch05, Tay01]. Matter [Lai01, ASRS08, Bat07, BCK^{+}03, BMW09, CN04, DK04, May01]. matter-antimatter [DK04]. Maxwell [MNV09]. mean [BHR03, FZ03, KSH^{+}06]. mean-field [BHR03, FZ03, KSH^{+}06]. measurement [BDV07, Sch04]. measurements [BUS08]. Measuring [FT03, Hal06]. mechanics [AB02, Sch04, Tay01, YR09]. mechanisms [Vio00]. media [AT06, BBF^{+}08, FIM05, PNHRD07, PNHRD09]. medium [HDB^{+}09, Mei09a, Mei09b]. melting [Str09]. melts [FFM^{+}03]. membranes [dJOS03]. mesh [Bec00]. mesoscopics [GHL^{+}06, GHL^{+}09]. Metallic [AKS01, FFM^{+}03, Hee01, KIC^{+}08]. metals [ABGH09, Gon03, Mac01, MvdL09, SRS05, Ste01, Ste06]. metamaterials [BBF^{+}08]. method [BCP08, WWAF06]. methods [BMH08]. Microfluidics [SQ05]. microgravity [BHLD07]. micromechanical [SHS06]. microscopes [HFS03]. Microscopic [SRS05]. microscopy [Gie03]. microtraps [FZ07]. Microwave [BCK^{+}03, Smo07]. Mills [SSJ04]. minimum [Gin04]. mirages [FH03]. Missile [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. mixing [GGN03, Kob09]. Mn [JSM^{+}06]. Mod [BFK^{+}04b, BFK^{+}04a, EHM14, GKMDA08b, GHL^{+}09, Gre00, HKP^{+}07a, JTLJ06a, KMN^{+}07b, Leg03, Mei09b, Nam10, PNHRD09, Zan01, ’t 02]. Mode [Das04]. Mode-coupling [Das04]. model [ABV^{+}05, CMPN^{+}05, FZ03, SBNC06, IN08, KO01]. Modeling [CFCT09, BDV07, RWP05]. models [BHR03, DPS04, Dyr06, PGT00]. Modern [EHM09a, Sah02, OS00]. modification [ABD^{+}06]. Molecular [FS00]. molecules [CSP09, JTLJ06a, JTLJ06b, KGJ06, SS03]. Momentum [PNHRD07, PNHRD09]. money [YR09]. Monte [FMNR01]. motion [CB09, Gre07, Sid09]. motors [HM09]. Mott [LNW06]. Multipolar [SCA^{+}09]. Muon [KO01].

nanoclusters [BF05]. nanofabrication [Ost05]. nanofluidics [SHR08]. nanoliter [SQ05]. nanomagnetism [Bad06]. nanoscale [HM09, Wie09]. nanostructures [SHB04]. nanotube [CCG^{+}01]. nanotubes [CBR07, LML02]. National [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. Nature [MvdL09]. Near [VP07]. Near-field [VP07]. Nematic [vTS00]. networks [AB02, DGM08]. neuroscience [RVSA06]. Neutrino [GGN03, AEE08, BGV02, Kos03]. neutrinos [AEE08, Dav03, KT01]. neutron [DHJ03]. NMR [Gre07, VC04]. Nobel [Abr04, Alf01, CW02, Dav03, Fer08, Gia03, Gin04, Gla06, Gro05, Grü08, Hal06, Hän06, Hee01, Ket02, Kob09, Kos03, Kro01, Leg04, Mac01, Mas09, Mat07, Nam09, Nam10, Pol05, Shi01, Smo07, Vel00, Wil05, ’t 00, ’t 02]. noise [SSGO07]. Non [NSS^{+}08, Ste01, Ste06]. Non-Abelian [NSS^{+}08]. Non-Fermi-liquid [Ste01, Ste06]. Noncommutative [DN01]. noncontact [VP07]. Nonequilibrium [EHM09b, Ódo04, WHG01, EHM14]. Nonlinear [MS06, SHS06, AVS09, BK00, BH07, BGK^{+}02]. Nonlocal [TBBH05]. Note [BFK^{+}04b, BFK^{+}04a, GKMDA08b, HKP^{+}07a, JTLJ06a, KMN^{+}07b, Mei09b, Nam10]. novel [ABD^{+}06, Mac01]. Nuclear [LMP03, BHR03, CMPN^{+}05, DHJ03, EHM09a, Gre98, Gre00, LPT03, PW07, SBNC06, WM09]. nuclei [DHJ03, Fra01]. nucleus [BDS08]. Numerical [BCP08].

Observation [KT01]. observational [Uza03]. Observations [NW00]. Odd [BVE05]. offsets [Kro01]. One [Gla06]. Onsager [ES06]. onset [Gro00]. Opportunities [Bad06, GHL^{+}06, GHL^{+}09]. Optical [LML02, BGK^{+}02, CT05, CY03, Hal06, KMN^{+}07a, KMN^{+}07b, LR09, MO06, Sah02]. Optics [CSP09, FIM05, MTB06, Bat07, BK00]. Optimal [WSG08]. Orbital [KK08]. Orbital-dependent [KK08]. Order [Sac03, SSGO07, TG09]. organic [GPM06, Mac01]. organized [SHB04]. Origin [DK04, Fer08, Wid02]. origins [OS00, Zur03]. oscillation [BGV02]. oscillations [MZVW01]. oscillator [SHS06]. other [Vio00]. our [Fer01]. oxides [DRS05].

Pairing [DHJ03, TK00, MM03]. paradigm [ABV^{+}05, SCA^{+}09, Wil05]. paradox [RDB^{+}09, Wil05]. parameter [BFE00]. particle [AP06, FG03, Hel01, Nam09, Nam10, dA07]. Particles [FGV01, Ugg05]. passage [KTS07]. Passion [Hän06]. Past [BHLD07]. Patterns [AT06]. Payment [Ano07]. perturbation [BdGDG01]. Phase [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c, BMW09, Kas08, Tso09, BKV05, GKI04, vLRVW07, Sac03]. phases [CCG^{+}01, JT02, MRMN03, Pfl09]. phenomena [AKS01, ABV^{+}05, BHLD07, BVE05, DGM08, LW05, RK02, SHR08]. Phonons [BdGDG01]. photoassociation [JTLJ06a, JTLJ06b]. photoemission [DHS03]. photoexcited [RK02]. photon [Luk03, MS06]. photon-photon [MS06]. photon-plasma [MS06]. Photonic [IS06, DGC06, KMN^{+}07a, KMN^{+}07b]. Photons [LW05, RBH01]. photoproduction [Kla02]. Phys [BFK^{+}04b, BFK^{+}04a, EHM14, GKMDA08b, GHL^{+}09, Gre00, HKP^{+}07a, JTLJ06a, KMN^{+}07b, Leg03, Mei09b, Nam10, PNHRD09, Zan01, ’t 02]. Physical [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c, Vio00, ZD08, Gin04, MT00, MT05, MTN08, PGT00]. Physically [RWP05]. physicists [Bec05]. Physics [Boo04, DRS05, ESL09, GHL^{+}06, GHL^{+}09, LNW06, Spr08b, Alf01, BDZ08, BGP^{+}09, CFL09, DRW06, DHZ06, FG03, GNS02, JL09, KI09, KO01, Lan09, LLLG04, MM03, MNV09, Mei09a, Mei09b, Nam09, Nam10, OW07, Pir04, RDM00, SJ01, SQ05, TKKTBW01, WM09]. physisorbed [BDV07]. picture [BM03]. pinches [RDR06, RDM00]. plane [SSJ04]. plane-wave [SSJ04]. plane-wave/super [SSJ04]. planets [BHLL01, Gon03]. plasma [ESL09, MS06, SE09, TG09]. plasma-based [ESL09]. plasmas [Boo04, GR09, MI09, Ost05, RWP05, Tho09, ZMD04]. Plasmonics [BBF^{+}08]. Podolsky [RDB^{+}09]. point [SRS05]. Policy [Ano07]. polyacetylene [Shi01]. polyelectrolytes [JL09, Vio00]. polymer [OW07]. polymeric [Hee01]. polymers [Hee01, Mac01, PLB08, Shi01, Str09, Tso09]. polynomial [WWAF06]. positron [Tho09]. power [NM02, RDR06]. practical [SBPC^{+}09]. precise [BUS08]. precision [Hän06]. premelted [DRW06]. Present [Hur03, BHLD07]. pressure [MRMN03]. Prices [Ano07]. primer [Kam02]. principle [Bou02]. principles [RVSA06]. Print [Ano07]. Prize [Mat07]. probe [Deu05, HFS03]. problem [Sch04]. problems [May01]. processes [LMP03]. Production [KGJ06]. Progress [BDV07]. properties [BF05, BdGDG01, CFCT09, CBR07, NGP^{+}09, SHB04]. protein [BM03, PGT00]. proton [Bas05, GF04, Wei03]. Proximity [Buz05]. pseudoscalar [GF04]. Publisher [BFK^{+}04b, BFK^{+}04a, GKMDA08b, HKP^{+}07a, JTLJ06a, KMN^{+}07b, Mei09b, Nam10]. Pulsed [WH00]. pulses [SS03].

Q [DW08]. QCD [CN04, Gro05]. quanta [Gla06]. Quantum [BvL05, DC08, FMNR01, GRTZ02, HHHH09, LBMW03, MJPH05, MSS01, PT04, SIGA05, Alh00, BM07, BRS07, BKV05, BCP08, Cha03, DPS04, EHM09b, EHM14, FH03, GMD02, Gre98, Gre00, HKP^{+}07a, HKP^{+}07b, JRFG04, KMN^{+}07a, KMN^{+}07b, vLRVW07, LR09, MZVW01, MS03, NSS^{+}08, RBH01, RM02, Sac03, SBPC^{+}09, Sch04, Sha08, Tay01, VC04, Ved02, Zur03, vdWDE^{+}03]. Quantum-state [MSS01, LR09]. quark [ASRS08, Wil00]. Quarkonia [EGMR08]. quarkonium [BPSV05]. quasi [DMW00]. quasi-two-dimensional [DMW00]. quasielastic [BDS08]. Quasielectric [Kro01]. qubits [KMN^{+}07a, KMN^{+}07b]. quest [ECS04].

radiation [Smo07]. radiative [VP07, vTS00]. Random [PW07, WM09, BBF^{+}08]. range [JTLJ06a, JTLJ06b]. rare [Hur03]. ray [DKI^{+}03, Gia03, GR09, KP07, KS01, Pir04, RA00, RRG01]. Rayleigh [AGL09]. rays [NW00]. reactions [JRFG04]. Reactive [Ost05]. Reactor [BGV02]. Reactor-based [BGV02]. Real [Bec00, KMM09]. Real-space [Bec00]. reappraisal [KF08]. recommended [MT00, MT05, MTN08]. reduce [RGC^{+}08]. reduction [PLB08]. Reference [BRS07]. reflected [Gre07]. reflection [Bee08]. reflections [Deu05]. refrigeration [GHL^{+}06, GHL^{+}09]. regime [MTB06]. reheating [BTW06]. related [AKS01, BdGDG01, BVE05, Hel01]. relation [GKI04]. relative [Ved02]. relativistic [MTB06, Ugg05]. Relativity [PT04]. renormalization [BCP08, Sch05, VG07]. Report [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. research [MI09, Spr08b, Zde02]. resistivity [GCH03]. resolved [DHS03]. resonance [Gre98, Gre00]. resonances [KGJ06]. Resonant [BGK^{+}02, KS01, BKM02]. resonators [BBF^{+}08]. Rev [BFK^{+}04b, BFK^{+}04a, EHM14, GKMDA08b, GHL^{+}09, Gre00, HKP^{+}07a, JTLJ06a, KMN^{+}07b, Leg03, Mei09b, Nam10, PNHRD09, Zan01, ’t 02]. Review [JBWS00]. RHIC [JBWS00]. Richardson [DPS04]. Role [SKC02, Mac01, Ved02]. roots [JO01]. Rosen [RDB^{+}09]. Rotating [Fet09, Fra01]. rules [BRS07]. RuO [MM03].

S1 [BFK^{+}04b, BFK^{+}04a]. Saint [Deu05]. Saint-James [Deu05]. Saturation [GCH03]. scale [AGL09, BKV05, HFS03, SQ05, Wie09]. Scaled [AP06]. scales [AP06, ZMD04]. Scaling [JMST04, GKMDA08a, GKMDA08b, Sid09]. Scanning [FKMA^{+}07, HFS03]. scattering [BDS08, DH07, GG08, GR09, JTLJ06a, JTLJ06b, KS01, dA07]. scenarios [JBWS00]. Scientific [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. searches [BCK^{+}03]. security [RGC^{+}08, SBPC^{+}09]. seen [Leg04]. Self [BHR03, Hel01, SHB04]. Self-consistent [BHR03]. self-driven [Hel01]. self-organized [SHB04]. Semiconducting [Hee01]. semiconductor [SHB04]. semiconductors [CT05, JSM^{+}06, RK02]. sequencing [ZD08]. shear [Gro00]. sheared [Ter00]. sheets [Wit07]. shell [CMPN^{+}05]. Si [Zan01, Zan00]. Simons [Mar05]. Simple [VG07, ABV^{+}05]. simulations [FMNR01, LML02, RG00, SKC02]. Single [BHL02, GPM06, LBMW03]. Single-bubble [BHL02]. single-crystal [GPM06]. size [GKMDA08a, GKMDA08b]. smectic [dJOS03]. Soap [Mor07]. social [CFL09]. Society [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. soft [Kam02]. solar [Dav03]. solids [DS00, FMNR01, KS01, PLT02, Sid09]. solitons [SY07]. solvable [DPS04]. Some [Leg01, Leg03, OW07, CW02, OS00]. sonoluminescence [BHL02]. space [Bec00, Kas08, ZMD04].

Spatiotemporal [SSGO07]. spectra [CT05, KS01, PW07]. spectroscopies [BKM02]. spectroscopy [FKMA^{+}07, JTLJ06a, JTLJ06b, Sid09]. speculative [JBWS00]. Spin [Wie09, Bas05, Fie07, Grü08, MM03]. spin-incoherent [Fie07]. spin-triplet [MM03]. Spins [HKP^{+}07b, HKP^{+}07a]. Spintronics [ZFS04, Fer08]. spiral [HDB^{+}09]. Spontaneous [Fra01, Nam09, Nam10]. spotlighting [Spr08b]. spreading [BEI^{+}09]. SR [SBK00, MM03]. stability [CM09]. standard [SBNC06, IN08, KO01]. Star [JL09, LK04]. Star-branched [JL09]. Stars [Gon03, DHJ03, LMP03, WHW02]. state [LR09, MSS01, SBK00, TRR00]. states [BVZ06, Luk03, MZVW01, MvdL09]. Statistical [AB02, CFL09, OW07, YR09, Alh00]. statistics [EHM09b, EHM14]. status [Hur03, Uza03]. Stress [Wit07]. String [BW06]. strings [IN08, Mar05]. stripes [KBF^{+}03]. strong [Lai01, SS03, Ugg05]. strongly [BMW09, Sha08]. Structural [BF05, SHB04]. Structure [JRFG04, KLLW07, SJ01, dJOS03, Bas05, BHR03, CMPN^{+}05, DW08, KIC^{+}08, KSH^{+}06, RA00, RM02, WM09]. structures [AVS09, BVE05]. studied [BKM02]. studies [DHS03, SBK00]. Study [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c]. Studying [Wil00]. substitutional [vdWC02]. super [SSJ04, BGP^{+}09]. Superconducting [Pfl09, CFCT09, Gre98, Gre00, JT02]. superconductivity [ASRS08, BVE05, CN04, DHZ04, Gin04, KW09, LNW06, MM03]. superconductor [BVE05, Buz05]. superconductor-ferromagnet [BVE05, Buz05]. superconductors [Abr04, ABGH09, BVZ06, BT05, DHS03, Deu05, FKMA^{+}07, HM02, KBF^{+}03, MZVW01, Sac03, SBK00, TK00]. Supercontinuum [DGC06]. supercooled [Das04, FFM^{+}03]. Superfluid [DP02, Leg04, GKMDA08a, GKMDA08b]. superfluidity [Gin04, WHG01]. superselection [BRS07]. supersonic [LK04]. Supersymmetric [SY07]. supersymmetry [SS00]. Suppression [Ter00]. surface [BAP05]. surfaces [TG09]. survey [Gre07]. symmetry [Cre01, Fra01, Nam09, Nam10, TK00]. synchronization [ABV^{+}05]. Synchrotrons [Wei03]. Synthetic [Mac01]. System [BFK^{+}04a]. Systems [BFK^{+}04b, BFK^{+}04c, AFOV08, BLVE07, BCP08, DMW00, DHJ03, DPS04, EHM09b, EHM14, GNS02, Hel01, Ódo04, RZ00, SCA^{+}09].

T [BT05, HM02, KW09]. tau [DHZ06]. teaching [Kro01]. Technical [BFK^{+}04b, BFK^{+}04a, BFK^{+}04c, NM02]. techniques [Bec00, VC04]. technology [Alf01, Sah02]. tell [Mas09]. temperature [FKMA^{+}07, KBF^{+}03, LNW06, PLT02, RWP05]. Tests [SBNC06]. Their [Smo07, EGMR08, JL09, Uza03]. theorem [SKC02]. theorems [EHM09b, EHM14]. theoretic [Tho09]. Theoretical [AT06, RA00, Uza03]. Theories [HFS03, BPSV05, MJPH05, MS03]. theorist [Leg04]. Theory [And04, BFE00, FH03, GPS08, JSM^{+}06, Kla02, KK08, PLB08, RK02, Alh00, AP06, BdGDG01, BM07, Bec00, BH07, BHLL01, CB09, Das04, DHZ04, DN01, EHM09a, ES06, FZ03, KMM09, KSH^{+}06, Mar05, OS00, PT04, RG00, TRR00, Tay01, Ved02]. thermal [PLT02]. thermodynamic [BF05]. thermodynamics [vdWC02]. thermometry [GHL^{+}06, GHL^{+}09]. thin [CM09, DRS05, DMW00]. thin-film [DRS05]. Thomson [GR09]. time [ZMD04]. tomography [LR09]. tool [Ost05]. top [Wil00]. topological [Mar05, NSS^{+}08]. topology [Kas08, OW07]. Tracing [Voi05]. Traffic [Hel01]. transfer [AGL09, BKM02, VP07, vTS00]. transistors [GPM06]. transition [Das04, DKI^{+}03, Dyr06, GKMDA08a, GKMDA08b]. transitions [BKV05, EGMR08, EM08, GK08, Kas08, vLRVW07, Sac03, Tso09]. transparency [FIM05]. Transport [SHR08, CBR07, GPM06, HM09, SJ01, Ter00, vdWDE^{+}03]. trapped [Fet09, LBMW03]. Trapping [Luk03]. trends [LPT03]. tricks [Kro01]. triplet [BVE05, MM03]. tube [BM03]. tunable [KGJ06]. tunneling [Bee08, FKMA^{+}07]. turbulence [ES06, FGV01, Gro00, LK04, PLB08, Ter00, ZMD04]. turbulent [AGL09]. tutorial [Bec05]. two [AKS01, DMW00, TRR00]. two-electron [TRR00]. type [MZVW01, SBK00, Abr04]. type-II [MZVW01, SBK00, Abr04].

Ultracold [JTLJ06b, BDZ08, FZ07, GPS08, JTLJ06a]. ultrafast [RK02]. ultrahigh [NW00]. ultrahigh-energy [NW00]. ultrarelativistic [Tho09]. unconventional [BVZ06, CFCT09]. Understanding [Sid09, FT03]. unified [CMPN^{+}05]. Universality [DS00, Ódo04]. universe [FT03, Hog00]. Unruh [CHM08]. Unusual [BBF^{+}08]. UPt [JT02]. using [BKM02]. utilization [Smo07].

V [JSM^{+}06, MRMN03]. values [MT00, MT05, MTN08]. vaporization [WH00]. variable [LR09]. variables [BvL05]. variation [Uza03]. versatile [Ost05]. versus [ORR02]. VI [MRMN03]. via [KGJ06]. vibrations [vdWC02]. view [CMPN^{+}05, SRS05]. violating [BFE00]. violation [IN08, Kob09, Mas09]. viscous [VG07]. Vortex [MZVW01, Abr04, AJ04, SBK00].

W [TKKTBW01]. Waiting [BUS08]. wall [PLB08]. wall-bounded [PLB08]. Wang [Spr08c]. warm [HDB^{+}09]. warming [RGC^{+}08]. wave [PNHRD07, PNHRD09]. wave/super [SSJ04]. waves [Grü08, Ket02]. weak [DP02, GF04, LMP03, Vel00]. weak-interaction [LMP03]. Weakly [Pos06, And04]. wealth [YR09]. Wei [Spr08c]. well [Gin04]. Wetting [BEI^{+}09]. Which [Spr08c]. within [RGC^{+}08]. works [Spr08a]. world [AK02].

X [DKI^{+}03, Gia03, GR09, KP07, KS01, RA00, RRG01]. X-ray [DKI^{+}03, Gia03, GR09, KP07, KS01, RA00, RRG01]. XXI [Gin04].

Yang [SSJ04]. years [CW02, Gla06].

References

Albert:2002:SMC

[AB02] Réka Albert and Albert-László Barabási. Statistical mechanics of complex networks. Reviews of Modern Physics, 74 (1):47–97, January 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.74.47;   http:/
/rmp.aps.org/abstract/RMP/v74/i1/p47_1.

Ahn:2006:EMN

[ABD^{+}06] C. H. Ahn, A. Bhattacharya, M. Di Ventra, J. N. Eckstein, C. Daniel Frisbie, M. E. Gershenson, A. M. Goldman, I. H. Inoue, J. Mannhart, Andrew J. Millis, Alberto F. Morpurgo, Douglas Natelson, and Jean-Marc Triscone. Electrostatic modification of novel materials. Reviews of Modern Physics, 78(4): 1185–1212, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1185;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1185_1.

Alloul:2009:DCM

[ABGH09] H. Alloul, J. Bobroff, M. Gabay, and P. J. Hirschfeld. Defects in correlated metals and superconductors. Reviews of Modern Physics, 81(1):45–108, January 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.45;   http://rmp.aps.org/abstract/RMP/v81/i1/p45_
1.

Abrikosov:2004:NLT

[Abr04] A. A. Abrikosov. Nobel Lecture: Type-II superconductors and the vortex lattice. Reviews of Modern Physics, 76(3):975–979, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.76.975;   http://rmp.aps.org/
abstract/RMP/v76/i3/p975_1.

Acebron:2005:KMS

[ABV^{+}05] Juan A. Acebrón, L. L. Bonilla, Conrad J. Pérez Vicente, Félix Ritort, and Renato Spigler. The Kuramoto model: a simple paradigm for synchronization phenomena. Reviews of Modern Physics, 77(1):137–185, January 2005. CODEN RMPHAT.

ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.77.137;
http://rmp.aps.org/abstract/RMP/v77/i1/p137_1.

Avignone:2008:DBD

[AEE08] Frank T. Avignone III, Steven R. Elliott, and Jonathan Engel. Double beta decay, Majorana neutrinos, and neutrino mass. Reviews of Modern Physics, 80(2):481–516, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.481;   http://rmp.aps.org/
abstract/RMP/v80/i2/p481_1.

Amico:2008:EMB

[AFOV08] Luigi Amico, Rosario Fazio, Andreas Osterloh, and Vlatko Vedral. Entanglement in many-body systems. Reviews of Modern Physics, 80(2):517–576, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.80.517;
http://rmp.aps.org/abstract/RMP/v80/i2/p517_1;   https:
//www.math.utah.edu/pub/bibnet/authors/f/fazio-riccardo.
bib.

Ahlers:2009:HTL

[AGL09] Guenter Ahlers, Siegfried Grossmann, and Detlef Lohse. Heat transfer and large scale dynamics in turbulent Rayleigh–Bénard convection. Reviews of Modern Physics, 81(2):503–537, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.81.503;   http://rmp.aps.org/
abstract/RMP/v81/i2/p503_1.

Altshuler:2004:CEV

[AJ04] E. Altshuler and T. H. Johansen. Colloquium: Experiments in vortex avalanches. Reviews of Modern Physics, 76(2):471– 487, April 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.76.471;   http://rmp.aps.
org/abstract/RMP/v76/i2/p471_1.

Aranson:2002:WCG

[AK02] Igor S. Aranson and Lorenz Kramer. The world of the complex Ginzburg–Landau equation. Reviews of Modern Physics, 74(1):99–143, January 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.74.99;   http:/
/rmp.aps.org/abstract/RMP/v74/i1/p99_1.

Abrahams:2001:MBR

[AKS01] Elihu Abrahams, Sergey V. Kravchenko, and Myriam P. Sarachik. Metallic behavior and related phenomena in two dimensions. Reviews of Modern Physics, 73(2):251–266, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.251;   http://rmp.aps.org/
abstract/RMP/v73/i2/p251_1.

Alferov:2001:NLD

[Alf01] Zhores I. Alferov. Nobel Lecture: The double heterostructure concept and its applications in physics, electronics, and technology. Reviews of Modern Physics, 73(3):767–782, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.767;   http://rmp.aps.org/
abstract/RMP/v73/i3/p767_1.

Alhassid:2000:STQ

[Alh00] Y. Alhassid. The statistical theory of quantum dots. Reviews of Modern Physics, 72(4):895–968, October 2000. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
72.895;   http://rmp.aps.org/abstract/RMP/v72/i4/p895_
1.

Andersen:2004:TWI

[And04] Jens O. Andersen. Theory of the weakly interacting Bose gas. Reviews of Modern Physics, 76(2):599–639, April 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.76.599;   http://rmp.aps.org/
abstract/RMP/v76/i2/p599_1.

Anonymous:2007:ANP

[Ano07] Anonymous. Announcement new prices and payment policy for color figures in the print journal. Reviews of Modern Physics, 79 (3):799, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.79.799;   http://rmp.aps.
org/abstract/RMP/v79/i3/p799_1.

Ashbaugh:2006:CSP

[AP06] Henry S. Ashbaugh and Lawrence R. Pratt. Colloquium: Scaled particle theory and the length scales of hydrophobicity. Reviews of Modern Physics, 78(1):159–178, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.159;   http://rmp.aps.org/
abstract/RMP/v78/i1/p159_1.

Alford:2008:CSD

[ASRS08] Mark G. Alford, Andreas Schmitt, Krishna Rajagopal, and Thomas Schäfer. Color superconductivity in dense quark matter. Reviews of Modern Physics, 80(4):1455–1515, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.80.1455;   http://rmp.aps.org/abstract/RMP/
v80/i4/p1455_1.

Aranson:2006:PCB

[AT06] Igor S. Aranson and Lev S. Tsimring. Patterns and collective behavior in granular media: Theoretical concepts. Reviews of Modern Physics, 78(2):641–692, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.78.641;
http://rmp.aps.org/abstract/RMP/v78/i2/p641_1.

Aguirre:2009:FSN

[AVS09] Jacobo Aguirre, Ricardo L. Viana, and Miguel A. F. Sanjuán. Fractal structures in nonlinear dynamics. Reviews of Modern Physics, 81(1):333–386, January 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.333;
http://rmp.aps.org/abstract/RMP/v81/i1/p333_1.

Bader:2006:CON

[Bad06] S. D. Bader. Colloquium: Opportunities in nanomagnetism. Reviews of Modern Physics, 78(1):1–15, January 2006. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.78.1;   http://rmp.aps.org/abstract/
RMP/v78/i1/p1_1.

Balibar:2005:SHC

[BAP05] Sébastien Balibar, Harry Alles, and Alexander Ya. Parshin. The surface of helium crystals. Reviews of Modern Physics, 77(1): 317–370, January 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.77.317;   http://rmp.
aps.org/abstract/RMP/v77/i1/p317_1.

Bass:2005:SSP

[Bas05] Steven D. Bass. The spin structure of the proton. Reviews of Modern Physics, 77(4):1257–1302, October 2005. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
77.1257;   http://rmp.aps.org/abstract/RMP/v77/i4/p1257_
1.

Batelaan:2007:CIK

[Bat07] H. Batelaan. Colloquium: Illuminating the Kapitza–Dirac effect with electron matter optics. Reviews of Modern Physics, 79(3):929–941, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.79.929;   http://rmp.
aps.org/abstract/RMP/v79/i3/p929_1.

Bliokh:2008:CUR

[BBF^{+}08] Konstantin Y. Bliokh, Yury P. Bliokh, Valentin Freilikher, Sergey Savel’ev, and Franco Nori. Colloquium: Unusual resonators: Plasmonics, metamaterials, and random media. Reviews of Modern Physics, 80(4):1201–1213, October 2008. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.80.1201;   http://rmp.aps.org/abstract/RMP/
v80/i4/p1201_1.

Bradley:2003:MCS

[BCK^{+}03] Richard Bradley, John Clarke, Darin Kinion, Leslie J. Rosenberg, Karl van Bibber, Seishi Matsuki, Michael Mück, and Pierre Sikivie. Microwave cavity searches for dark-matter axions. Reviews of Modern Physics, 75(3):777–817, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.777;   http://rmp.aps.org/
abstract/RMP/v75/i3/p777_1.

Bulla:2008:NRG

[BCP08] Ralf Bulla, Theo A. Costi, and Thomas Pruschke. Numerical renormalization group method for quantum impurity systems. Reviews of Modern Physics, 80(2):395–450, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.395;   http://rmp.aps.org/
abstract/RMP/v80/i2/p395_1.

Baroni:2001:PRC

[BdGDG01] Stefano Baroni, Stefano de Gironcoli, Andrea Dal Corso, and Paolo Giannozzi. Phonons and related crystal properties from density-functional perturbation theory. Reviews of Modern Physics, 73(2):515–562, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.73.515;
http://rmp.aps.org/abstract/RMP/v73/i2/p515_1.

Benhar:2008:IQE

[BDS08] Omar Benhar, Donal Day, and Ingo Sick. Inclusive quasielastic electron-nucleus scattering. Reviews of Modern Physics, 80(1): 189–224, January 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.80.189;   http://rmp.
aps.org/abstract/RMP/v80/i1/p189_1.

Bruch:2007:PMM

[BDV07] L. W. Bruch, R. D. Diehl, and J. A. Venables. Progress in the measurement and modeling of physisorbed layers. Reviews of Modern Physics, 79(4):1381–1454, October 2007. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
79.1381;   http://rmp.aps.org/abstract/RMP/v79/i4/p1381_
1.

Bloch:2008:MBP

[BDZ08] Immanuel Bloch, Jean Dalibard, and Wilhelm Zwerger. Manybody physics with ultracold gases. Reviews of Modern Physics, 80(3):885–964, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.80.885;   http://rmp.
aps.org/abstract/RMP/v80/i3/p885_1.

Beck:2000:RSM

[Bec00] Thomas L. Beck. Real-space mesh techniques in densityfunctional theory. Reviews of Modern Physics, 72(4):1041– 1080, October 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.72.1041;   http:
//rmp.aps.org/abstract/RMP/v72/i4/p1041_1.

Bechhoefer:2005:FPT

[Bec05] John Bechhoefer. Feedback for physicists: a tutorial essay on control. Reviews of Modern Physics, 77(3):783–836, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.783;   http://rmp.aps.org/
abstract/RMP/v77/i3/p783_1.

Beenakker:2008:CAR

[Bee08] C. W. J. Beenakker. Colloquium: Andreev reflection and Klein tunneling in graphene. Reviews of Modern Physics, 80(4): 1337–1354, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1337;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1337_1.

Bonn:2009:WS

[BEI^{+}09] Daniel Bonn, Jens Eggers, Joseph Indekeu, Jacques Meunier, and Etienne Rolley. Wetting and spreading. Reviews of Modern Physics, 81(2):739–805, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.739;
http://rmp.aps.org/abstract/RMP/v81/i2/p739_1.

Baletto:2005:SPN

[BF05] Francesca Baletto and Riccardo Ferrando. Structural properties of nanoclusters: Energetic, thermodynamic, and kinetic effects. Reviews of Modern Physics, 77(1):371–423, January 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.371;   http://rmp.aps.org/
abstract/RMP/v77/i1/p371_1.

Bertolini:2000:TCV

[BFE00] Stefano Bertolini, Marco Fabbrichesi, and Jan O. Eeg. Theory of the CP-violating parameter ϵ′/ϵ. Reviews of Modern Physics, 72(1):65–93, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.72.65;   http:/
/rmp.aps.org/abstract/RMP/v72/i1/p65_1.

Barton:2004:PNRb

[BFK^{+}04a] David K. Barton, Roger Falcone, Daniel Kleppner, Frederick K. Lamb, Ming K. Lau, Harvey L. Lynch, David Moncton, David Montague, David E. Mosher, William Priedhorsky, Maury Tigner, and David R. Vaughan. Publisher’s note: Report of the American Physical Society Study Group on Boost-Phase Intercept System for National Missile Defense: Scientific and technical issues [Rev. Mod. Phys. 76, S1 (2004)]. Reviews of Modern Physics, 76(4):1307– 1322, October 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.76.1307;   http:
//rmp.aps.org/abstract/RMP/v76/i4/p1307_1.

Barton:2004:PNRa

[BFK^{+}04b] David K. Barton, Roger Falcone, Daniel Kleppner, Frederick K. Lamb, Ming K. Lau, Harvey L. Lynch, David Moncton, David Montague, David E. Mosher, William Priedhorsky, Maury Tigner, and David R. Vaughan. Publisher’s note: Report of the American Physical Society Study Group on Boost-Phase Intercept Systems for National Missile Defense: Scientific and Technical Issues [Rev. Mod. Phys. 76, S1 (2004)]. Reviews of Modern Physics, 76(3):S425– 425, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.76.S425;   http:
//rmp.aps.org/abstract/RMP/v76/i3/pS425_1.

Barton:2004:RAP

[BFK^{+}04c] David K. Barton, Roger Falcone, Daniel Kleppner, Frederick K. Lamb, Ming K. Lau, Harvey L. Lynch, David Moncton, David Montague, David E. Mosher, William Priedhorsky, Maury Tigner, and David R. Vaughan. Report of the American Physical Society Study Group on Boost-Phase Intercept Systems for National Missile Defense: Scientific and technical issues. Reviews of Modern Physics, 76(3):S1–S424, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.76.S1;   http://rmp.aps.org/abstract/RMP/
v76/i3/pS1_1.

Budker:2002:RNM

[BGK^{+}02] D. Budker, W. Gawlik, D. F. Kimball, S. M. Rochester, V. V. Yashchuk, and A. Weis. Resonant nonlinear magnetooptical effects in atoms. Reviews of Modern Physics, 74(4): 1153–1201, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.74.1153;   http:
//rmp.aps.org/abstract/RMP/v74/i4/p1153_1.

Browder:2009:NPS

[BGP^{+}09] Thomas E. Browder, Tim Gershon, Dan Pirjol, Amarjit Soni, and Jure Zupan. New physics at a Super Flavor Factory. Reviews of Modern Physics, 81(4):1887–1941, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.81.1887;   http://rmp.aps.org/abstract/RMP/
v81/i4/p1887_1.

Bemporad:2002:RBN

[BGV02] Carlo Bemporad, Giorgio Gratta, and Petr Vogel. Reactor-based neutrino oscillation experiments. Reviews of Modern Physics, 74 (2):297–328, April 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.74.297;   http://rmp.
aps.org/abstract/RMP/v74/i2/p297_1.

Brizard:2007:FNG

[BH07] A. J. Brizard and T. S. Hahm. Foundations of nonlinear gyrokinetic theory. Reviews of Modern Physics, 79(2):421–468, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.79.421;   http://rmp.aps.org/
abstract/RMP/v79/i2/p421_1.

Brenner:2002:SBS

[BHL02] Michael P. Brenner, Sascha Hilgenfeldt, and Detlef Lohse. Single-bubble sonoluminescence. Reviews of Modern Physics, 74(2):425–484, April 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.74.425;   http://rmp.
aps.org/abstract/RMP/v74/i2/p425_1.

Barmatz:2007:CPM

[BHLD07] M. Barmatz, Inseob Hahn, J. A. Lipa, and R. V. Duncan. Critical phenomena in microgravity: Past, present, and future. Reviews of Modern Physics, 79(1):1–52, January 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.79.1;   http://rmp.aps.org/abstract/
RMP/v79/i1/p1_1.

Burrows:2001:TBD

[BHLL01] Adam Burrows, W. B. Hubbard, J. I. Lunine, and James Liebert. The theory of brown dwarfs and extrasolar giant planets. Reviews of Modern Physics, 73(3):719–765, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.719;   http://rmp.aps.org/
abstract/RMP/v73/i3/p719_1.

Bender:2003:SCM

[BHR03] Michael Bender, Paul-Henri Heenen, and Paul-Gerhard Reinhard. Self-consistent mean-field models for nuclear structure. Reviews of Modern Physics, 75(1):121–180, January 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.121;   http://rmp.aps.org/
abstract/RMP/v75/i1/p121_1.

Brabec:2000:IFC

[BK00] Thomas Brabec and Ferenc Krausz. Intense few-cycle laser fields: Frontiers of nonlinear optics. Reviews of Modern Physics, 72(2):545–591, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.72.545;   http://rmp.
aps.org/abstract/RMP/v72/i2/p545_1.

Bruhwiler:2002:CTD

[BKM02] P. A. Brühwiler, O. Karis, and N. Mårtensson. Charge-transfer dynamics studied using resonant core spectroscopies. Reviews of Modern Physics, 74(3):703–740, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.74.703;
http://rmp.aps.org/abstract/RMP/v74/i3/p703_1.

Belitz:2005:HGS

[BKV05] D. Belitz, T. R. Kirkpatrick, and Thomas Vojta. How generic scale invariance influences quantum and classical phase transitions. Reviews of Modern Physics, 77(2):579–632, April 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.579;   http://rmp.aps.org/
abstract/RMP/v77/i2/p579_1.

Beloborodov:2007:GES

[BLVE07] I. S. Beloborodov, A. V. Lopatin, V. M. Vinokur, and K. B. Efetov. Granular electronic systems. Reviews of Modern Physics, 79(2):469–518, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.79.469;   http://rmp.
aps.org/abstract/RMP/v79/i2/p469_1.

Banavar:2003:CGA

[BM03] Jayanth R. Banavar and Amos Maritan. Colloquium: Geometrical approach to protein folding: a tube picture. Reviews of Modern Physics, 75(1):23–34, January 2003. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
75.23;   http://rmp.aps.org/abstract/RMP/v75/i1/p23_
1.

Bartlett:2007:CCT

[BM07] Rodney J. Bartlett and Monika Musial. Coupled-cluster theory in quantum chemistry. Reviews of Modern Physics, 79(1): 291–352, January 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.79.291;   http://rmp.
aps.org/abstract/RMP/v79/i1/p291_1.

Belkic:2008:FBM

[BMH08] Dzevad Belkić, Ivan Mancev, and Jocelyn Hanssen. Four-body methods for high-energy ion-atom collisions. Reviews of Modern Physics, 80(1):249–314, January 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.80.249;
http://rmp.aps.org/abstract/RMP/v80/i1/p249_1.

Braun-Munzinger:2009:CPD

[BMW09] P. Braun-Munzinger and J. Wambach. Colloquium: Phase diagram of strongly interacting matter. Reviews of Modern Physics, 81(3):1031–1050, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1031;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1031_1.

Boozer:2004:PMC

[Boo04] Allen H. Boozer. Physics of magnetically confined plasmas. Reviews of Modern Physics, 76(4):1071–1141, October 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.76.1071;   http://rmp.aps.org/abstract/RMP/
v76/i4/p1071_1.

Bousso:2002:HP

[Bou02] Raphael Bousso. The holographic principle. Reviews of Modern Physics, 74(3):825–874, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.74.825;
http://rmp.aps.org/abstract/RMP/v74/i3/p825_1.

Brambilla:2005:EFT

[BPSV05] Nora Brambilla, Antonio Pineda, Joan Soto, and Antonio Vairo. Effective-field theories for heavy quarkonium. Reviews of Modern Physics, 77(4):1423–1496, October 2005. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
77.1423;   http://rmp.aps.org/abstract/RMP/v77/i4/p1423_
1.

Bartlett:2007:RFS

[BRS07] Stephen D. Bartlett, Terry Rudolph, and Robert W. Spekkens. Reference frames, superselection rules, and quantum information. Reviews of Modern Physics, 79(2):555–609, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.79.555;   http://rmp.aps.org/
abstract/RMP/v79/i2/p555_1.

Basov:2005:EHS

[BT05] D. N. Basov and T. Timusk. Electrodynamics of high-T_{c} superconductors. Reviews of Modern Physics, 77(2):721–779, April 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.721;   http://rmp.aps.org/
abstract/RMP/v77/i2/p721_1.

Bassett:2006:IDR

[BTW06] Bruce A. Bassett, Shinji Tsujikawa, and David Wands. Inflation dynamics and reheating. Reviews of Modern Physics, 78 (2):537–589, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.78.537;   http://rmp.
aps.org/abstract/RMP/v78/i2/p537_1.

Buras:2008:WPM

[BUS08] Andrzej J. Buras, Selma Uhlig, and Felix Schwab. Waiting for precise measurements of K^{+} → π^{+}νν̄ and K_{L} → π^{0}νν̄. Reviews of Modern Physics, 80(3):965–1007, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.965;   http://rmp.aps.org/
abstract/RMP/v80/i3/p965_1.

Buzdin:2005:PES

[Buz05] A. I. Buzdin. Proximity effects in superconductor-ferromagnet heterostructures. Reviews of Modern Physics, 77(3):935–976, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.935;   http://rmp.aps.org/
abstract/RMP/v77/i3/p935_1.

Bergeret:2005:OTS

[BVE05] F. S. Bergeret, A. F. Volkov, and K. B. Efetov. Odd triplet superconductivity and related phenomena in superconductorferromagnet structures. Reviews of Modern Physics, 77(4):

1321–1373, October 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1321;   http:
//rmp.aps.org/abstract/RMP/v77/i4/p1321_1.

Braunstein:2005:QIC

[BvL05] Samuel L. Braunstein and Peter van Loock. Quantum information with continuous variables. Reviews of Modern Physics, 77 (2):513–577, April 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.77.513;   http://rmp.
aps.org/abstract/RMP/v77/i2/p513_1.

Balatsky:2006:IIS

[BVZ06] A. V. Balatsky, I. Vekhter, and Jian-Xin Zhu. Impurityinduced states in conventional and unconventional superconductors. Reviews of Modern Physics, 78(2):373–433, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.373;   http://rmp.aps.org/
abstract/RMP/v78/i2/p373_1.

Battefeld:2006:SGC

[BW06] Thorsten Battefeld and Scott Watson. String gas cosmology. Reviews of Modern Physics, 78(2):435–454, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.435;   http://rmp.aps.org/
abstract/RMP/v78/i2/p435_1.

Cary:2009:HTG

[CB09] John R. Cary and Alain J. Brizard. Hamiltonian theory of guiding-center motion. Reviews of Modern Physics, 81(2):693– 738, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.81.693;   http://rmp.aps.
org/abstract/RMP/v81/i2/p693_1.

Charlier:2007:ETP

[CBR07] Jean-Christophe Charlier, Xavier Blase, and Stephan Roche. Electronic and transport properties of nanotubes. Reviews of Modern Physics, 79(2):677–732, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.677;
http://rmp.aps.org/abstract/RMP/v79/i2/p677_1.

Calbi:2001:CPG

[CCG^{+}01] M. Mercedes Calbi, Milton W. Cole, Silvina M. Gatica, Mary J. Bojan, and George Stan. Condensed phases of gases inside nanotube bundles. Reviews of Modern Physics, 73(4):857–865, October 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.73.857;   http://rmp.aps.
org/abstract/RMP/v73/i4/p857_1.

Christensen-Dalsgaard:2002:H

[CD02] Jørgen Christensen-Dalsgaard. Helioseismology. Reviews of Modern Physics, 74(4):1073–1129, October 2002. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
74.1073;   http://rmp.aps.org/abstract/RMP/v74/i4/p1073_
1.

Capone:2009:CMU

[CFCT09] Massimo Capone, Michele Fabrizio, Claudio Castellani, and Erio Tosatti. Colloquium: Modeling the unconventional superconducting properties of expanded A_{3}C_{60} fullerides. Reviews of Modern Physics, 81(2):943–958, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.943;
http://rmp.aps.org/abstract/RMP/v81/i2/p943_1.

Castellano:2009:SPS

[CFL09] Claudio Castellano, Santo Fortunato, and Vittorio Loreto. Statistical physics of social dynamics. Reviews of Modern Physics, 81(2):591–646, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.81.591;   http://rmp.
aps.org/abstract/RMP/v81/i2/p591_1.

Chang:2003:CLL

[Cha03] A. M. Chang. Chiral Luttinger liquids at the fractional quantum Hall edge. Reviews of Modern Physics, 75(4):1449– 1505, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1449;   http:
//rmp.aps.org/abstract/RMP/v75/i4/p1449_1.

Crispino:2008:UEA

[CHM08] Luís C. B. Crispino, Atsushi Higuchi, and George E. A. Matsas. The Unruh effect and its applications. Reviews of Modern Physics, 80(3):787–838, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.80.787;
http://rmp.aps.org/abstract/RMP/v80/i3/p787_1.

Chu:2004:ECM

[Chu04] K. R. Chu. The electron cyclotron maser. Reviews of Modern Physics, 76(2):489–540, April 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.76.489;
http://rmp.aps.org/abstract/RMP/v76/i2/p489_1.

Craster:2009:DST

[CM09] R. V. Craster and O. K. Matar. Dynamics and stability of thin liquid films. Reviews of Modern Physics, 81(3): 1131–1198, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1131;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1131_1.

Caurier:2005:SMU

[CMPN^{+}05] E. Caurier, G. Martínez-Pinedo, F. Nowacki, A. Poves, and A. P. Zuker. The shell model as a unified view of nuclear structure. Reviews of Modern Physics, 77(2):427–488, April 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.427;   http://rmp.aps.org/
abstract/RMP/v77/i2/p427_1.

Casalbuoni:2004:ISC

[CN04] Roberto Casalbuoni and Giuseppe Nardulli. Inhomogeneous superconductivity in condensed matter and QCD. Reviews of Modern Physics, 76(1):263–320, January 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.76.263;
http://rmp.aps.org/abstract/RMP/v76/i1/p263_1.

Creutz:2001:ACS

[Cre01] Michael Creutz. Aspects of chiral symmetry and the lattice. Reviews of Modern Physics, 73(1):119–150, January 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.119;   http://rmp.aps.org/
abstract/RMP/v73/i1/p119_1.

Cronin:2009:OIA

[CSP09] Alexander D. Cronin, Jörg Schmiedmayer, and David E. Pritchard. Optics and interferometry with atoms and molecules. Reviews of Modern Physics, 81(3):1051–1129, July 2009. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.81.1051;   http://rmp.aps.org/abstract/RMP/
v81/i3/p1051_1.

Cardona:2005:IEO

[CT05] Manuel Cardona and M. L. W. Thewalt. Isotope effects on the optical spectra of semiconductors. Reviews of Modern Physics, 77(4):1173–1224, October 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1173;   http:
//rmp.aps.org/abstract/RMP/v77/i4/p1173_1.

Cornell:2002:NLB

[CW02] E. A. Cornell and C. E. Wieman. Nobel Lecture: Bose–Einstein condensation in a dilute gas, the first 70 years and some recent experiments. Reviews of Modern Physics, 74(3):875–893, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.875;   http://rmp.aps.org/
abstract/RMP/v74/i3/p875_1.

Cundiff:2003:CFO

[CY03] Steven T. Cundiff and Jun Ye. Colloquium: Femtosecond optical frequency combs. Reviews of Modern Physics, 75(1):325–342, January 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.75.325;   http://rmp.aps.
org/abstract/RMP/v75/i1/p325_1.

deAbajo:2007:CLS

[dA07] F. J. García de Abajo. Colloquium: Light scattering by particle and hole arrays. Reviews of Modern Physics, 79(4): 1267–1290, October 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.79.1267;   http:
//rmp.aps.org/abstract/RMP/v79/i4/p1267_1.

Das:2004:MCT

[Das04] Shankar P. Das. Mode-coupling theory and the glass transition in supercooled liquids. Reviews of Modern Physics, 76 (3):785–851, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.76.785;   http://rmp.
aps.org/abstract/RMP/v76/i3/p785_1.

Davis:2003:NLH

[Dav03] Raymond Davis, Jr. Nobel Lecture: a half-century with solar neutrinos. Reviews of Modern Physics, 75(3):985–994, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.985;   http://rmp.aps.org/
abstract/RMP/v75/i3/p985_1.

Das:2008:CQA

[DC08] Arnab Das and Bikas K. Chakrabarti. Colloquium: Quantum annealing and analog quantum computation. Reviews of Modern Physics, 80(3):1061–1081, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.80.
1061;   http://rmp.aps.org/abstract/RMP/v80/i3/p1061_
1.

Deutscher:2005:ASJ

[Deu05] Guy Deutscher. Andreev–Saint-James reflections: a probe of cuprate superconductors. Reviews of Modern Physics, 77(1): 109–135, January 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.77.109;   http://rmp.
aps.org/abstract/RMP/v77/i1/p109_1.

Dudley:2006:SGP

[DGC06] John M. Dudley, Goëry Genty, and Stéphane Coen. Supercontinuum generation in photonic crystal fiber. Reviews of Modern Physics, 78(4):1135–1184, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.78.
1135;   http://rmp.aps.org/abstract/RMP/v78/i4/p1135_
1.

Dorogovtsev:2008:CPC

[DGM08] S. N. Dorogovtsev, A. V. Goltsev, and J. F. F. Mendes. Critical phenomena in complex networks. Reviews of Modern Physics, 80 (4):1275–1335, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1275;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1275_1.

Devereaux:2007:ILS

[DH07] Thomas P. Devereaux and Rudi Hackl. Inelastic light scattering from correlated electrons. Reviews of Modern Physics, 79(1): 175–233, January 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.79.175;   http://rmp.
aps.org/abstract/RMP/v79/i1/p175_1.

Dean:2003:PNS

[DHJ03] D. J. Dean and M. Hjorth-Jensen. Pairing in nuclear systems: from neutron stars to finite nuclei. Reviews of Modern Physics, 75(2):607–656, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.75.607;   http://rmp.
aps.org/abstract/RMP/v75/i2/p607_1.

Damascelli:2003:ARP

[DHS03] Andrea Damascelli, Zahid Hussain, and Zhi-Xun Shen. Angleresolved photoemission studies of the cuprate superconductors. Reviews of Modern Physics, 75(2):473–541, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.473;   http://rmp.aps.org/
abstract/RMP/v75/i2/p473_1.

Demler:2004:TAS

[DHZ04] Eugene Demler, Werner Hanke, and Shou-Cheng Zhang. SO(5) theory of antiferromagnetism and superconductivity. Reviews of Modern Physics, 76(3):909–974, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.76.909;
http://rmp.aps.org/abstract/RMP/v76/i3/p909_1.

Davier:2006:PHT

[DHZ06] Michel Davier, Andreas Höcker, and Zhiqing Zhang. The physics of hadronic tau decays. Reviews of Modern Physics, 78(4): 1043–1109, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1043;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1043_1.

deJeu:2003:SFS

[dJOS03] Wim H. de Jeu, Boris I. Ostrovskii, and Arcadi N. Shalaginov. Structure and fluctuations of smectic membranes. Reviews of Modern Physics, 75(1):181–235, January 2003. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
75.181;   http://rmp.aps.org/abstract/RMP/v75/i1/p181_
1.

Dine:2004:OMA

[DK04] Michael Dine and Alexander Kusenko. Origin of the matterantimatter asymmetry. Reviews of Modern Physics, 76(1): 1–30, January 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.76.1;   http://
rmp.aps.org/abstract/RMP/v76/i1/p1_1.

Douglas:2007:FC

[DK07] Michael R. Douglas and Shamit Kachru. Flux compactification. Reviews of Modern Physics, 79(2):733–796, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.79.733;   http://rmp.aps.org/
abstract/RMP/v79/i2/p733_1.

Deslattes:2003:XRT

[DKI^{+}03] Richard D. Deslattes, Ernest G. Kessler, Jr., P. Indelicato, L. de Billy, E. Lindroth, and J. Anton. X-ray transition energies: new approach to a comprehensive evaluation. Reviews of Modern Physics, 75(1):35–99, January 2003. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
75.35;   http://rmp.aps.org/abstract/RMP/v75/i1/p35_
1.

DeBell:2000:DEM

[DMW00] K. De’Bell, A. B. MacIsaac, and J. P. Whitehead. Dipolar effects in magnetic thin films and quasi-two-dimensional systems. Reviews of Modern Physics, 72(1):225–257, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.72.225;   http://rmp.aps.org/
abstract/RMP/v72/i1/p225_1.

Douglas:2001:NFT

[DN01] Michael R. Douglas and Nikita A. Nekrasov. Noncommutative field theory. Reviews of Modern Physics, 73(4):977–1029, October 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.73.977;   http://rmp.aps.
org/abstract/RMP/v73/i4/p977_1.

Davis:2002:SJW

[DP02] J. C. Davis and R. E. Packard. Superfluid ^{3}He Josephson weak links. Reviews of Modern Physics, 74(3):741–773, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.741;   http://rmp.aps.org/
abstract/RMP/v74/i3/p741_1.

Dukelsky:2004:CES

[DPS04] J. Dukelsky, S. Pittel, and G. Sierra. Colloquium: Exactly solvable Richardson–Gaudin models for many-body quantum systems. Reviews of Modern Physics, 76(3):643–662, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.76.643;   http://rmp.aps.org/
abstract/RMP/v76/i3/p643_1.

Dawber:2005:PTF

[DRS05] M. Dawber, K. M. Rabe, and J. F. Scott. Physics of thinfilm ferroelectric oxides. Reviews of Modern Physics, 77(4): 1083–1130, October 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1083;   http:
//rmp.aps.org/abstract/RMP/v77/i4/p1083_1.

Dash:2006:PPI

[DRW06] J. G. Dash, A. W. Rempel, and J. S. Wettlaufer. The physics of premelted ice and its geophysical consequences. Reviews of Modern Physics, 78(3):695–741, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.78.695;
http://rmp.aps.org/abstract/RMP/v78/i3/p695_1.

Dyre:2000:UAC

[DS00] Jeppe C. Dyre and Thomas B. Schrøder. Universality of ac conduction in disordered solids. Reviews of Modern Physics, 72 (3):873–892, July 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.72.873;   http://rmp.
aps.org/abstract/RMP/v72/i3/p873_1.

Drechsel:2008:HSL

[DW08] Dieter Drechsel and Thomas Walcher. Hadron structure at low Q^{2}. Reviews of Modern Physics, 80(3):731–785, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.731;   http://rmp.aps.org/
abstract/RMP/v80/i3/p731_1.

Dyre:2006:CGT

[Dyr06] Jeppe C. Dyre. Colloquium: The glass transition and elastic models of glass-forming liquids. Reviews of Modern Physics, 78 (3):953–972, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.78.953;   http://rmp.
aps.org/abstract/RMP/v78/i3/p953_1.

Endres:2004:CQH

[ECS04] R. G. Endres, D. L. Cox, and R. R. P. Singh. Colloquium: The quest for high-conductance DNA. Reviews of Modern Physics, 76(1):195–214, January 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.76.195;   http:
//rmp.aps.org/abstract/RMP/v76/i1/p195_1.

Eichten:2008:QTT

[EGMR08] Estia Eichten, Stephen Godfrey, Hanna Mahlke, and Jonathan L. Rosner. Quarkonia and their transitions. Reviews of Modern Physics, 80(3):1161–1193, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.80.
1161;   http://rmp.aps.org/abstract/RMP/v80/i3/p1161_
1.

Epelbaum:2009:MTN

[EHM09a] E. Epelbaum, H.-W. Hammer, and Ulf-G. Meißner. Modern theory of nuclear forces. Reviews of Modern Physics, 81(4): 1773–1825, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1773;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1773_1.

Esposito:2009:NFF

[EHM09b] Massimiliano Esposito, Upendra Harbola, and Shaul Mukamel. Nonequilibrium fluctuations, fluctuation theorems, and counting statistics in quantum systems. Reviews of Modern Physics, 81(4):1665–1702, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1665;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1665_1. See erratum

[EHM14].

Esposito:2014:ENF

[EHM14] Massimiliano Esposito, Upendra Harbola, and Shaul Mukamel. Erratum: Nonequilibrium fluctuations, fluctuation theorems, and counting statistics in quantum systems [Rev. Mod. Phys. 81, 1665 (2009)]. Reviews of Modern Physics, 86(??):1125–??, March 2014. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://journals.aps.org/

rmp/abstract/10.1103/RevModPhys.86.1125. See [EHM09b].

Evers:2008:AT

[EM08] Ferdinand Evers and Alexander D. Mirlin. Anderson transitions. Reviews of Modern Physics, 80(4):1355–1417, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.80.1355;   http://rmp.aps.org/abstract/RMP/
v80/i4/p1355_1.

Eyink:2006:OTH

[ES06] Gregory L. Eyink and Katepalli R. Sreenivasan. Onsager and the theory of hydrodynamic turbulence. Reviews of Modern Physics, 78(1):87–135, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.78.87;   http:/
/rmp.aps.org/abstract/RMP/v78/i1/p87_1.

Esarey:2009:PLD

[ESL09] E. Esarey, C. B. Schroeder, and W. P. Leemans. Physics of laser-driven plasma-based electron accelerators. Reviews of Modern Physics, 81(3):1229–1285, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.81.
1229;   http://rmp.aps.org/abstract/RMP/v81/i3/p1229_
1.

Ferriere:2001:IEO

[Fer01] Katia M. Ferrière. The interstellar environment of our galaxy. Reviews of Modern Physics, 73(4):1031–1066, October 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.73.1031;   http://rmp.aps.org/abstract/RMP/
v73/i4/p1031_1.

Fert:2008:NLO

[Fer08] Albert Fert. Nobel Lecture: Origin, development, and future of spintronics. Reviews of Modern Physics, 80(4):1517– 1530, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1517;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1517_1.

Fetter:2009:RTB

[Fet09] Alexander L. Fetter. Rotating trapped Bose–Einstein condensates. Reviews of Modern Physics, 81(2):647–691, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.81.647;   http://rmp.aps.org/
abstract/RMP/v81/i2/p647_1.

Faupel:2003:DMG

[FFM^{+}03] Franz Faupel, Werner Frank, Michael-Peter Macht, Helmut Mehrer, Volkmar Naundorf, Klaus Rätzke, Herbert R. Schober, Suman K. Sharma, and Helmar Teichler. Diffusion in metallic glasses and supercooled melts. Reviews of Modern Physics, 75 (1):237–280, January 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.75.237;   http://rmp.
aps.org/abstract/RMP/v75/i1/p237_1.

Fabjan:2003:CPP

[FG03] Christian W. Fabjan and Fabiola Gianotti. Calorimetry for particle physics. Reviews of Modern Physics, 75(4):1243– 1286, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1243;   http:
//rmp.aps.org/abstract/RMP/v75/i4/p1243_1.

Falkovich:2001:PFF

[FGV01] G. Falkovich, K. Gawȩdzki, and M. Vergassola. Particles and fields in fluid turbulence. Reviews of Modern Physics, 73(4): 913–975, October 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.73.913;   http://rmp.
aps.org/abstract/RMP/v73/i4/p913_1.

Fiete:2003:CTQ

[FH03] Gregory A. Fiete and Eric J. Heller. Colloquium: Theory of quantum corrals and quantum mirages. Reviews of Modern Physics, 75(3):933–948, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.75.933;
http://rmp.aps.org/abstract/RMP/v75/i3/p933_1.

Fiete:2007:CSI

[Fie07] Gregory A. Fiete. Colloquium: The spin-incoherent Luttinger liquid. Reviews of Modern Physics, 79(3):801–820, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.79.801;   http://rmp.aps.org/
abstract/RMP/v79/i3/p801_1.

Fleischhauer:2005:EIT

[FIM05] Michael Fleischhauer, Atac Imamoglu, and Jonathan P. Marangos. Electromagnetically induced transparency: Optics in coherent media. Reviews of Modern Physics, 77(2):633–673, April 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.633;   http://rmp.aps.org/
abstract/RMP/v77/i2/p633_1.

Fischer:2007:STS

[FKMA^{+}07] Øystein Fischer, Martin Kugler, Ivan Maggio-Aprile, Christophe Berthod, and Christoph Renner. Scanning tunneling spectroscopy of high-temperature superconductors. Reviews of Modern Physics, 79(1):353–419, January 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.353;
http://rmp.aps.org/abstract/RMP/v79/i1/p353_1.

Foulkes:2001:QMC

[FMNR01] W. M. C. Foulkes, L. Mitas, R. J. Needs, and G. Rajagopal. Quantum Monte Carlo simulations of solids. Reviews of Modern Physics, 73(1):33–83, January 2001. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
73.33;   http://rmp.aps.org/abstract/RMP/v73/i1/p33_
1.

Frauendorf:2001:SSB

[Fra01] Stefan Frauendorf. Spontaneous symmetry breaking in rotating nuclei. Reviews of Modern Physics, 73(2):463–514, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.463;   http://rmp.aps.org/
abstract/RMP/v73/i2/p463_1.

Feldmeier:2000:MDF

[FS00] Hans Feldmeier and Jürgen Schnack. Molecular dynamics for fermions. Reviews of Modern Physics, 72(3):655–688, July 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.72.655;   http://rmp.aps.org/
abstract/RMP/v72/i3/p655_1.

Freedman:2003:CMU

[FT03] Wendy L. Freedman and Michael S. Turner. Colloquium: Measuring and understanding the universe. Reviews of Modern Physics, 75(4):1433–1447, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.75.
1433;   http://rmp.aps.org/abstract/RMP/v75/i4/p1433_
1.

Freericks:2003:EDM

[FZ03] J. K. Freericks and V. Zlatić. Exact dynamical mean-field theory of the Falicov–Kimball model. Reviews of Modern Physics, 75 (4):1333–1382, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1333;   http:
//rmp.aps.org/abstract/RMP/v75/i4/p1333_1.

Fortagh:2007:MMU

[FZ07] József Fortágh and Claus Zimmermann. Magnetic microtraps for ultracold atoms. Reviews of Modern Physics, 79(1):235–289, January 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.79.235;   http://rmp.aps.
org/abstract/RMP/v79/i1/p235_1.

Gunnarsson:2003:CSE

[GCH03] O. Gunnarsson, M. Calandra, and J. E. Han. Colloquium: Saturation of electrical resistivity. Reviews of Modern Physics, 75 (4):1085–1099, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1085;   http:
//rmp.aps.org/abstract/RMP/v75/i4/p1085_1.

Gorringe:2004:IPC

[GF04] Tim Gorringe and Harold W. Fearing. Induced pseudoscalar coupling of the proton weak interaction. Reviews of Modern Physics, 76(1):31–91, January 2004. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
76.31;   http://rmp.aps.org/abstract/RMP/v76/i1/p31_
1.

Gargioni:2008:ESA

[GG08] E. Gargioni and B. Grosswendt. Electron scattering from argon: Data evaluation and consistency. Reviews of Modern Physics, 80(2):451–480, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.80.451;   http://rmp.
aps.org/abstract/RMP/v80/i2/p451_1.

Gonzalez-Garcia:2003:NMM

[GGN03] M. C. Gonzalez-Garcia and Yosef Nir. Neutrino masses and mixing: evidence and implications. Reviews of Modern Physics, 75 (2):345–402, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.75.345;   http://rmp.
aps.org/abstract/RMP/v75/i2/p345_1.

Giazotto:2006:OMT

[GHL^{+}06] Francesco Giazotto, Tero T. Heikkilä, Arttu Luukanen, Alexander M. Savin, and Jukka P. Pekola. Opportunities for mesoscopics in thermometry and refrigeration: Physics and applications. Reviews of Modern Physics, 78(1):217–274, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.217;   http://rmp.aps.org/
abstract/RMP/v78/i1/p217_1. See erratum [GHL^{+}09].

Giazotto:2009:EOM

[GHL^{+}09] Francesco Giazotto, Tero T. Heikkilä, Arttu Luukanen, Alexander M. Savin, and Jukka P. Pekola. Erratum: Opportunities for mesoscopics in thermometry and refrigeration: Physics and applications [Rev. Mod. Phys. 78, 217 (2006)]. Reviews of Modern Physics, 81(3):1351, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1351;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1351_1. See [GHL^{+}06].

Giacconi:2003:NLD

[Gia03] Riccardo Giacconi. Nobel Lecture: The dawn of X-ray astronomy. Reviews of Modern Physics, 75(3):995–1010, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.995;   http://rmp.aps.org/
abstract/RMP/v75/i3/p995_1.

Giessibl:2003:AAF

[Gie03] Franz J. Giessibl. Advances in atomic force microscopy. Reviews of Modern Physics, 75(3):949–983, July 2003. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
75.949;   http://rmp.aps.org/abstract/RMP/v75/i3/p949_
1.

Ginzburg:2004:NLS

[Gin04] Vitaly L. Ginzburg. Nobel Lecture: On superconductivity and superfluidity (what I have and have not managed to do) as well as on the “physical minimum” at the beginning of the XXI century. Reviews of Modern Physics, 76(3):981–998, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.76.981;   http://rmp.aps.org/
abstract/RMP/v76/i3/p981_1.

Gwyn:2008:CGT

[GK08] Rhiannon Gwyn and Anke Knauf. Conifolds and geometric transitions. Reviews of Modern Physics, 80(4):1419– 1453, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1419;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1419_1.

Golubov:2004:CPR

[GKI04] A. A. Golubov, M. Yu. Kupriyanov, and E. Il’ichev. The current-phase relation in Josephson junctions. Reviews of Modern Physics, 76(2):411–469, April 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.76.411;
http://rmp.aps.org/abstract/RMP/v76/i2/p411_1.

Gasparini:2008:FSS

[GKMDA08a] Francis M. Gasparini, Mark O. Kimball, Kevin P. Mooney, and Manuel Diaz-Avila. Finite-size scaling of ^{4}He at the superfluid transition. Reviews of Modern Physics, 80(3): 1009–1059, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1009;   http:
//rmp.aps.org/abstract/RMP/v80/i3/p1009_1. See note

[GKMDA08b].

Gasparini:2008:PNF

[GKMDA08b] Francis M. Gasparini, Mark O. Kimball, Kevin P. Mooney, and Manuel Diaz-Avila. Publisher’s note: Finite-size scaling of ^{4}He at the superfluid transition [rev. mod. phys. 80, 1009 (2008)]. Reviews of Modern Physics, 80(3): 1195, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1195;   http:
//rmp.aps.org/abstract/RMP/v80/i3/p1195_1. See [GKMDA08a].

Glauber:2006:NLO

[Gla06] Roy J. Glauber. Nobel Lecture: One hundred years of light quanta. Reviews of Modern Physics, 78(4):1267– 1278, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1267;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1267_1.

Gailitis:2002:CLE

[GLP^{+}02] Agris Gailitis, Olgerts Lielausis, Ernests Platacis, Gunter Gerbeth, and Frank Stefani. Colloquium: Laboratory experiments on hydromagnetic dynamos. Reviews of Modern Physics, 74(4): 973–990, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.74.973;   http://rmp.
aps.org/abstract/RMP/v74/i4/p973_1.

Galindo:2002:ICC

[GMD02] A. Galindo and M. A. Martín-Delgado. Information and computation: Classical and quantum aspects. Reviews of Modern Physics, 74(2):347–423, April 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.74.347;
http://rmp.aps.org/abstract/RMP/v74/i2/p347_1.

Grosberg:2002:CPC

[GNS02] A. Yu. Grosberg, T. T. Nguyen, and B. I. Shklovskii. Colloquium: The physics of charge inversion in chemical and biological systems. Reviews of Modern Physics, 74(2):329–345, April 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.329;   http://rmp.aps.org/
abstract/RMP/v74/i2/p329_1.

Gonzalez:2003:CSP

[Gon03] Guillermo Gonzalez. Colloquium: Stars, planets, and metals. Reviews of Modern Physics, 75(1):101–120, January 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.101;   http://rmp.aps.org/
abstract/RMP/v75/i1/p101_1.

Gershenson:2006:CET

[GPM06] M. E. Gershenson, V. Podzorov, and A. F. Morpurgo. Colloquium: Electronic transport in single-crystal organic transistors. Reviews of Modern Physics, 78(3):973–989, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.973;   http://rmp.aps.org/
abstract/RMP/v78/i3/p973_1.

Giorgini:2008:TUA

[GPS08] Stefano Giorgini, Lev P. Pitaevskii, and Sandro Stringari. Theory of ultracold atomic Fermi gases. Reviews of Modern Physics, 80(4):1215–1274, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1215;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1215_1.

Glenzer:2009:XRT

[GR09] Siegfried H. Glenzer and Ronald Redmer. X-ray Thomson scattering in high energy density plasmas. Reviews of Modern Physics, 81(4):1625–1663, October 2009. CODEN RMPHAT.

ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.81.
1625;   http://rmp.aps.org/abstract/RMP/v81/i4/p1625_
1.

Greenberg:1998:ASQ

[Gre98] Ya. S. Greenberg. Application of superconducting quantum interference devices to nuclear magnetic resonance. Reviews of Modern Physics, 70(1):175–222, January 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
70.175;   http://rmp.aps.org/abstract/RMP/v70/i1/p175_

1. See erratum [Gre00].

Greenberg:2000:EAS

[Gre00] Ya. S. Greenberg. Erratum: Application of superconducting quantum interference devices to nuclear magnetic resonance [Rev. Mod. Phys. 70, 175 (1998)]. Reviews of Modern Physics, 72(1):329, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.72.329;   http://rmp.
aps.org/abstract/RMP/v72/i1/p329_1. See [Gre98].

Grebenkov:2007:NSR

[Gre07]  Denis S. Grebenkov.  NMR survey of reflected Brow-

nian motion. Reviews of Modern Physics, 79(3):1077– 1137, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.79.1077;   http:
//rmp.aps.org/abstract/RMP/v79/i3/p1077_1.

Grossmann:2000:OSF

[Gro00]  Siegfried Grossmann.  The onset of shear flow turbu-

lence. Reviews of Modern Physics, 72(2):603–618, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.72.603;   http://rmp.aps.org/
abstract/RMP/v72/i2/p603_1.

Gross:2005:NLD

[Gro05] David J. Gross. Nobel Lecture: The discovery of asymptotic freedom and the emergence of QCD. Reviews of Modern Physics, 77(3):837–849, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.77.837;   http://rmp.
aps.org/abstract/RMP/v77/i3/p837_1.

Gisin:2002:QC

[GRTZ02] Nicolas Gisin, Grégoire Ribordy, Wolfgang Tittel, and Hugo Zbinden. Quantum cryptography. Reviews of Modern Physics, 74(1):145–195, January 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.74.145;   http:
//rmp.aps.org/abstract/RMP/v74/i1/p145_1.

Grunberg:2008:NLS

[Grü08] Peter A. Grünberg. Nobel Lecture: From spin waves to giant magnetoresistance and beyond. Reviews of Modern Physics, 80 (4):1531–1540, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1531;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1531_1.

Hall:2006:NLD

[Hal06] John L. Hall. Nobel Lecture: Defining and measuring optical frequencies. Reviews of Modern Physics, 78(4):1279– 1295, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1279;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1279_1.

Hansch:2006:NLP

[Hän06] Theodor W. Hänsch. Nobel Lecture: Passion for precision. Reviews of Modern Physics, 78(4):1297–1309, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.78.1297;   http://rmp.aps.org/abstract/RMP/
v78/i4/p1297_1.

Haffner:2009:WIM

[HDB^{+}09] L. M. Haffner, R.-J. Dettmar, J. E. Beckman, K. Wood, J. D. Slavin, C. Giammanco, G. J. Madsen, A. Zurita, and R. J. Reynolds. The warm ionized medium in spiral galaxies. Reviews of Modern Physics, 81(3):969–997, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.81.969;   http://rmp.aps.org/
abstract/RMP/v81/i3/p969_1.

Heeger:2001:NLS

[Hee01] Alan J. Heeger. Nobel Lecture: Semiconducting and metallic polymers: The fourth generation of polymeric materials. Reviews of Modern Physics, 73(3):681–700, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.681;   http://rmp.aps.org/
abstract/RMP/v73/i3/p681_1.

Helbing:2001:TRS

[Hel01] Dirk Helbing. Traffic and related self-driven many-particle systems. Reviews of Modern Physics, 73(4):1067–1141, October 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.73.1067;   http://rmp.aps.org/abstract/
RMP/v73/i4/p1067_1.

Hofer:2003:TSP

[HFS03] Werner A. Hofer, Adam S. Foster, and Alexander L. Shluger. Theories of scanning probe microscopes at the atomic scale. Reviews of Modern Physics, 75(4):1287–1331, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.75.1287;   http://rmp.aps.org/abstract/RMP/
v75/i4/p1287_1.

Horodecki:2009:QE

[HHHH09] Ryszard Horodecki, Pawel Horodecki, Michal Horodecki, and Karol Horodecki. Quantum entanglement. Reviews of Modern Physics, 81(2):865–942, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.865;
http://rmp.aps.org/abstract/RMP/v81/i2/p865_1.

Hanson:2007:PNS

[HKP^{+}07a] R. Hanson, L. P. Kouwenhoven, J. R. Petta, S. Tarucha, and L. M. K. Vandersypen. Publisher’s note: Spins in few-electron quantum dots [Rev. Mod. Phys. 79, 1217 (2007)]. Reviews of Modern Physics, 79(4):1455, October 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.79.
1455;   http://rmp.aps.org/abstract/RMP/v79/i4/p1455_

1. See [HKP^{+}07b].

Hanson:2007:SFE

[HKP^{+}07b] R. Hanson, L. P. Kouwenhoven, J. R. Petta, S. Tarucha, and L. M. K. Vandersypen. Spins in few-electron quantum dots. Reviews of Modern Physics, 79(4):1217–1265, October 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.79.1217;   http://rmp.aps.org/abstract/RMP/

v79/i4/p1217_1. See note [HKP^{+}07a].

Hofmann:2000:DHE

[HM00] S. Hofmann and G. Münzenberg. The discovery of the heaviest elements. Reviews of Modern Physics, 72(3):733–767, July 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.72.733;   http://rmp.aps.org/
abstract/RMP/v72/i3/p733_1.

Hilgenkamp:2002:GBH

[HM02] H. Hilgenkamp and J. Mannhart. Grain boundaries in high-T_{c} superconductors. Reviews of Modern Physics, 74(2):485– 549, April 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.74.485;   http://rmp.aps.
org/abstract/RMP/v74/i2/p485_1.

Hanggi:2009:ABM

[HM09] Peter Hänggi and Fabio Marchesoni. Artificial Brownian motors: Controlling transport on the nanoscale. Reviews of Modern Physics, 81(1):387–442, January 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.387;
http://rmp.aps.org/abstract/RMP/v81/i1/p387_1.

Hogan:2000:WUJ

[Hog00] Craig J. Hogan. Why the universe is just so. Reviews of Modern Physics, 72(4):1149–1161, October 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.72.

1149;   http://rmp.aps.org/abstract/RMP/v72/i4/p1149_
1.

Hurth:2003:PSI

[Hur03] Tobias Hurth. Present status of inclusive rare B decays. Reviews of Modern Physics, 75(4):1159–1199, October 2003. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.75.1159;   http://rmp.aps.org/abstract/RMP/
v75/i4/p1159_1.

Ibrahim:2008:CVS

[IN08] Tarek Ibrahim and Pran Nath. CP violation from the Standard Model to strings. Reviews of Modern Physics, 80(2):577– 631, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.80.577;   http://rmp.aps.
org/abstract/RMP/v80/i2/p577_1.

Istrate:2006:PCH

[IS06] Emanuel Istrate and Edward H. Sargent. Photonic crystal heterostructures and interfaces. Reviews of Modern Physics, 78 (2):455–481, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.78.455;   http://rmp.
aps.org/abstract/RMP/v78/i2/p455_1.

Jaffe:2000:RSD

[JBWS00] R. L. Jaffe, W. Busza, F. Wilczek, and J. Sandweiss. Review of speculative “disaster scenarios” at RHIC. Reviews of Modern Physics, 72(4):1125–1140, October 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.72.
1125;   http://rmp.aps.org/abstract/RMP/v72/i4/p1125_
1.

Jusufi:2009:CSB

[JL09] Arben Jusufi and Christos N. Likos. Colloquium: Starbranched polyelectrolytes: The physics of their conformations and interactions. Reviews of Modern Physics, 81(4): 1753–1772, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1753;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1753_1.

Jones:2004:SLD

[JMST04] Bernard J. T. Jones, Vicent J. Martínez, Enn Saar, and Virginia Trimble. Scaling laws in the distribution of galaxies. Reviews of Modern Physics, 76(4):1211–1266, October 2004. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
76.1211;   http://rmp.aps.org/abstract/RMP/v76/i4/p1211_
1.

Jackson:2001:HRG

[JO01] J. D. Jackson and L. B. Okun. Historical roots of gauge invariance. Reviews of Modern Physics, 73(3):663–680, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.663;   http://rmp.aps.org/
abstract/RMP/v73/i3/p663_1.

Jensen:2004:SRQ

[JRFG04] A. S. Jensen, K. Riisager, D. V. Fedorov, and E. Garrido. Structure and reactions of quantum halos. Reviews of Modern Physics, 76(1):215–261, January 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.76.215;
http://rmp.aps.org/abstract/RMP/v76/i1/p215_1.

Jungwirth:2006:TFI

[JSM^{+}06] T. Jungwirth, Jairo Sinova, J. Masek, J. Kucera, and A. H. MacDonald. Theory of ferromagnetic (III,mn)V semiconductors. Reviews of Modern Physics, 78(3):809–864, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.809;   http://rmp.aps.org/
abstract/RMP/v78/i3/p809_1.

Joynt:2002:SPU

[JT02] Robert Joynt and Louis Taillefer. The superconducting phases of UPt_{3}. Reviews of Modern Physics, 74(1):235–294, January 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.235;   http://rmp.aps.org/
abstract/RMP/v74/i1/p235_1.

Jones:2006:PNU

[JTLJ06a] Kevin M. Jones, Eite Tiesinga, Paul D. Lett, and Paul S. Julienne. Publisher’s note: Ultracold photoassociation spectroscopy: Long-range molecules and atomic scattering [Rev. Mod. Phys. 78, 483 (2006)]. Reviews of Modern Physics, 78 (3):1041, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1041;   http:
//rmp.aps.org/abstract/RMP/v78/i3/p1041_1. See [JTLJ06b].

Jones:2006:UPS

[JTLJ06b] Kevin M. Jones, Eite Tiesinga, Paul D. Lett, and Paul S. Julienne. Ultracold photoassociation spectroscopy: Long-range molecules and atomic scattering. Reviews of Modern Physics, 78(2):483–535, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.78.483;   http://rmp.
aps.org/abstract/RMP/v78/i2/p483_1. See note [JTLJ06a].

Kamien:2002:GSM

[Kam02] Randall D. Kamien. The geometry of soft materials: a primer. Reviews of Modern Physics, 74(4):953–971, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.953;   http://rmp.aps.org/
abstract/RMP/v74/i4/p953_1.

Kastner:2008:PTC

[Kas08] Michael Kastner. Phase transitions and configuration space topology. Reviews of Modern Physics, 80(1):167–187, January 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.167;   http://rmp.aps.org/
abstract/RMP/v80/i1/p167_1.

Kivelson:2003:HDF

[KBF^{+}03] S. A. Kivelson, I. P. Bindloss, E. Fradkin, V. Oganesyan, J. M. Tranquada, A. Kapitulnik, and C. Howald. How to detect fluctuating stripes in the high-temperature superconductors. Reviews of Modern Physics, 75(4):1201– 1241, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1201;   http:
//rmp.aps.org/abstract/RMP/v75/i4/p1201_1.

Ketterle:2002:NLW

[Ket02] Wolfgang Ketterle. Nobel Lecture: When atoms behave as waves: Bose–Einstein condensation and the atom laser. Reviews of Modern Physics, 74(4):1131–1151, October 2002. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.74.1131;   http://rmp.aps.org/abstract/RMP/
v74/i4/p1131_1.

Kleman:2008:DDC

[KF08] M. Kleman and J. Friedel. Disclinations, dislocations, and continuous defects: a reappraisal. Reviews of Modern Physics, 80 (1):61–115, January 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.80.61;   http:/
/rmp.aps.org/abstract/RMP/v80/i1/p61_1.

Kohler:2006:PCM

[KGJ06] Thorsten Köhler, Krzysztof Góral, and Paul S. Julienne. Production of cold molecules via magnetically tunable Feshbach resonances. Reviews of Modern Physics, 78(4):1311– 1361, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1311;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1311_1.

Krausz:2009:AP

[KI09] Ferenc Krausz and Misha Ivanov. Attosecond physics. Reviews of Modern Physics, 81(1):163–234, January 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.163;   http://rmp.aps.org/abstract/RMP/v81/i1/p163_
1.

Katsnelson:2008:HMF

[KIC^{+}08] M. I. Katsnelson, V. Yu. Irkhin, L. Chioncel, A. I. Lichtenstein, and R. A. de Groot. Half-metallic ferromagnets: From band structure to many-body effects. Reviews of Modern Physics, 80 (2):315–378, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.80.315;   http://rmp.
aps.org/abstract/RMP/v80/i2/p315_1.

Kummel:2008:ODD

[KK08] Stephan Kümmel and Leeor Kronik. Orbital-dependent density functionals: Theory and applications. Reviews of Modern Physics, 80(1):3–60, January 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.80.3;
 http://rmp.aps.org/abstract/RMP/v80/i1/p3_1.

Klasen:2002:THP

[Kla02] Michael Klasen. Theory of hard photoproduction. Reviews of Modern Physics, 74(4):1221–1282, October 2002. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
74.1221;   http://rmp.aps.org/abstract/RMP/v74/i4/p1221_
1.

Kornyshev:2007:SIB

[KLLW07] Alexei A. Kornyshev, Dominic J. Lee, Sergey Leikin, and Aaron Wynveen. Structure and interactions of biological helices. Reviews of Modern Physics, 79(3):943–996, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.79.943;   http://rmp.aps.org/
abstract/RMP/v79/i3/p943_1.

Krechetnikov:2007:DII

[KM07] R. Krechetnikov and J. E. Marsden. Dissipation-induced instabilities in finite dimensions. Reviews of Modern Physics, 79 (2):519–553, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.79.519;   http://rmp.
aps.org/abstract/RMP/v79/i2/p519_1.

Klimchitskaya:2009:CFB

[KMM09] G. L. Klimchitskaya, U. Mohideen, and V. M. Mostepanenko. The Casimir force between real materials: Experiment and theory. Reviews of Modern Physics, 81(4):1827– 1885, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1827;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1827_1.

Kok:2007:LOQ

[KMN^{+}07a] Pieter Kok, W. J. Munro, Kae Nemoto, T. C. Ralph, Jonathan P. Dowling, and G. J. Milburn. Linear optical quantum computing with photonic qubits. Reviews of Modern Physics, 79(1):135–174, January 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.135;
http://rmp.aps.org/abstract/RMP/v79/i1/p135_1.  See

note [KMN^{+}07b].

Kok:2007:PNL

[KMN^{+}07b] Pieter Kok, W. J. Munro, Kae Nemoto, T. C. Ralph, Jonathan P. Dowling, and G. J. Milburn. Publisher’s note: Linear optical quantum computing with photonic qubits [Rev. Mod. Phys. 79, 135 (2007)]. Reviews of Modern Physics, 79(2): 797, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.79.797;   http://rmp.aps.
org/abstract/RMP/v79/i2/p797_1. See [KMN^{+}07a].

Kuno:2001:MDP

[KO01] Yoshitaka Kuno and Yasuhiro Okada. Muon decay and physics beyond the Standard Model. Reviews of Modern Physics, 73(1): 151–202, January 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.73.151;   http://rmp.
aps.org/abstract/RMP/v73/i1/p151_1.

Kobayashi:2009:NLC

[Kob09] Makoto Kobayashi. Nobel Lecture: CP violation and flavor mixing. Reviews of Modern Physics, 81(3):1019– 1025, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1019;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1019_1.

Koshiba:2003:NLB

[Kos03] Masatoshi Koshiba. Nobel Lecture: Birth of neutrino astrophysics. Reviews of Modern Physics, 75(3):1011–1020, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.75.1011;   http://rmp.aps.org/abstract/
RMP/v75/i3/p1011_1.

Kallman:2007:ADX

[KP07] T. R. Kallman and P. Palmeri. Atomic data for Xray astrophysics. Reviews of Modern Physics, 79(1):79– 133, January 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.79.79;   http://
rmp.aps.org/abstract/RMP/v79/i1/p79_1.

Kroemer:2001:NLQ

[Kro01] Herbert Kroemer. Nobel Lecture: Quasielectric fields and band offsets: teaching electrons new tricks. Reviews of Modern Physics, 73(3):783–793, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.73.783;
http://rmp.aps.org/abstract/RMP/v73/i3/p783_1.

Kotani:2001:RIX

[KS01] Akio Kotani and Shik Shin. Resonant inelastic X-ray scattering spectra for electrons in solids. Reviews of Modern Physics, 73(1):203–246, January 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.73.203;   http:
//rmp.aps.org/abstract/RMP/v73/i1/p203_1.

Kotliar:2006:ESC

[KSH^{+}06] G. Kotliar, S. Y. Savrasov, K. Haule, V. S. Oudovenko, O. Parcollet, and C. A. Marianetti. Electronic structure calculations with dynamical mean-field theory. Reviews of Modern Physics, 78(3):865–951, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.78.865;   http://rmp.
aps.org/abstract/RMP/v78/i3/p865_1.

Kajita:2001:OAN

[KT01] Takaaki Kajita and Yoji Totsuka. Observation of atmospheric neutrinos. Reviews of Modern Physics, 73(1):85– 118, January 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.73.85;   http://
rmp.aps.org/abstract/RMP/v73/i1/p85_1.

Kral:2007:CCC

[KTS07] Petr Král, Ioannis Thanopulos, and Moshe Shapiro. Colloquium: Coherently controlled adiabatic passage. Reviews of Modern Physics, 79(1):53–77, January 2007. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
79.53;   http://rmp.aps.org/abstract/RMP/v79/i1/p53_
1.

Kresin:2009:CEL

[KW09] V. Z. Kresin and S. A. Wolf. Colloquium: Electron-lattice interaction and its impact on high T_{c} superconductivity. Reviews of Modern Physics, 81(2):481–501, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.481;
http://rmp.aps.org/abstract/RMP/v81/i2/p481_1.

Lai:2001:MSM

[Lai01] Dong Lai. Matter in strong magnetic fields. Reviews of Modern Physics, 73(3):629–662, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.73.629;
http://rmp.aps.org/abstract/RMP/v73/i3/p629_1.

Langacker:2009:PHG

[Lan09] Paul Langacker. The physics of heavy Z′ gauge bosons. Reviews of Modern Physics, 81(3):1199–1228, July 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.1199;   http://rmp.aps.org/abstract/RMP/v81/i3/p1199_
1.

Leibfried:2003:QDS

[LBMW03] D. Leibfried, R. Blatt, C. Monroe, and D. Wineland. Quantum dynamics of single trapped ions. Reviews of Modern Physics, 75(1):281–324, January 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.75.281;   http:
//rmp.aps.org/abstract/RMP/v75/i1/p281_1.

Leggett:2001:BEC

[Leg01] Anthony J. Leggett. Bose–Einstein condensation in the alkali gases: Some fundamental concepts. Reviews of Modern Physics, 73(2):307–356, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.73.307;   http:
//rmp.aps.org/abstract/RMP/v73/i2/p307_1. See erratum

[Leg03].

Leggett:2003:EBE

[Leg03] Anthony J. Leggett. Erratum: Bose–Einstein condensation in the alkali gases: Some fundamental concepts [Rev. Mod. Phys. 73, 307 (2001)]. Reviews of Modern Physics, 75 (3):1083, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1083;   http:
//rmp.aps.org/abstract/RMP/v75/i3/p1083_1. See [Leg01].

Leggett:2004:NLS

[Leg04] Anthony J. Leggett. Nobel Lecture: Superfluid ^{3}He: the early days as seen by a theorist. Reviews of Modern Physics, 76 (3):999–1011, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.76.999;   http://rmp.
aps.org/abstract/RMP/v76/i3/p999_1.

Low:2004:CSF

[LK04] Mordecai-Mark Mac Low and Ralf S. Klessen. Control of star formation by supersonic turbulence. Reviews of Modern Physics, 76(1):125–194, January 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.76.125;   http:
//rmp.aps.org/abstract/RMP/v76/i1/p125_1.

Lister:2004:PDL

[LLLG04] G. G. Lister, J. E. Lawler, W. P. Lapatovich, and V. A. Godyak. The physics of discharge lamps. Reviews of Modern Physics, 76 (2):541–598, April 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.76.541;   http://rmp.
aps.org/abstract/RMP/v76/i2/p541_1.

Lucas:2002:OSE

[LML02] A. A. Lucas, F. Moreau, and Ph. Lambin. Optical simulations of electron diffraction by carbon nanotubes. Reviews of Modern Physics, 74(1):1–10, January 2002. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
74.1;   http://rmp.aps.org/abstract/RMP/v74/i1/p1_1.

Langanke:2003:NWI

[LMP03] K. Langanke and G. Martínez-Pinedo. Nuclear weak-interaction processes in stars. Reviews of Modern Physics, 75(3):819–862, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.819;   http://rmp.aps.org/
abstract/RMP/v75/i3/p819_1.

Lee:2006:DMI

[LNW06] Patrick A. Lee, Naoto Nagaosa, and Xiao-Gang Wen. Doping a Mott insulator: Physics of high-temperature superconductivity. Reviews of Modern Physics, 78(1):17–85, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.78.17;   http://rmp.aps.org/abstract/RMP/
v78/i1/p17_1.

Lunney:2003:RTD

[LPT03] D. Lunney, J. M. Pearson, and C. Thibault. Recent trends in the determination of nuclear masses. Reviews of Modern Physics, 75(3):1021–1082, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1021;   http:
//rmp.aps.org/abstract/RMP/v75/i3/p1021_1.

Lvovsky:2009:CVO

[LR09] A. I. Lvovsky and M. G. Raymer. Continuous-variable optical quantum-state tomography. Reviews of Modern Physics, 81(1): 299–332, January 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.81.299;   http://rmp.
aps.org/abstract/RMP/v81/i1/p299_1.

Lukin:2003:CTM

[Luk03] M. D. Lukin. Colloquium: Trapping and manipulating photon states in atomic ensembles. Reviews of Modern Physics, 75 (2):457–472, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.75.457;   http://rmp.
aps.org/abstract/RMP/v75/i2/p457_1.

Levin:2005:CPE

[LW05] Michael Levin and Xiao-Gang Wen. Colloquium: Photons and electrons as emergent phenomena. Reviews of Modern Physics, 77(3):871–879, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.77.871;   http://rmp.
aps.org/abstract/RMP/v77/i3/p871_1.

MacDiarmid:2001:NLS

[Mac01] Alan G. MacDiarmid. Nobel Lecture: “synthetic metals”: a novel role for organic polymers. Reviews of Modern Physics, 73 (3):701–712, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.73.701;   http://rmp.
aps.org/abstract/RMP/v73/i3/p701_1.

Marino:2005:CST

[Mar05]  Marcos Mariño.  Chern–Simons theory and topological

strings. Reviews of Modern Physics, 77(2):675–720, April 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.675;   http://rmp.aps.org/
abstract/RMP/v77/i2/p675_1.

Maskawa:2009:NLW

[Mas09] Toshihide Maskawa. Nobel Lecture: What does CP violation tell us? Reviews of Modern Physics, 81(3):1027– 1030, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1027;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1027_1.

Mather:2007:NLB

[Mat07] John C. Mather. Nobel Lecture: From the Big Bang to the Nobel Prize and beyond. Reviews of Modern Physics, 79(4):

1331–1348, October 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.79.1331;   http:
//rmp.aps.org/abstract/RMP/v79/i4/p1331_1.

Maynard:2001:AAC

[May01] J. D. Maynard. Acoustical analogs of condensed-matter problems. Reviews of Modern Physics, 73(2):401–417, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.401;   http://rmp.aps.org/
abstract/RMP/v73/i2/p401_1.

Meiksin:2009:PIM

[Mei09a] Avery A. Meiksin. The physics of the intergalactic medium. Reviews of Modern Physics, 81(4):1405–1469, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.81.1405;   http://rmp.aps.org/abstract/RMP/

v81/i4/p1405_1. See note [Mei09b].

Meiksin:2009:PNP

[Mei09b] Avery A. Meiksin. Publisher’s note: The physics of the intergalactic medium [Rev. Mod. Phys. 81, 1405 (2009)]. Reviews of Modern Physics, 81(4):1943, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.81.
1943;   http://rmp.aps.org/abstract/RMP/v81/i4/p1943_

1. See [Mei09a].

Morfill:2009:CPI

[MI09] Gregor E. Morfill and Alexei V. Ivlev. Complex plasmas: An interdisciplinary research field. Reviews of Modern Physics, 81 (4):1353–1404, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1353;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1353_1.

Maier:2005:QCT

[MJPH05] Thomas Maier, Mark Jarrell, Thomas Pruschke, and Matthias H. Hettler. Quantum cluster theories. Reviews of Modern Physics, 77(3):1027–1080, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1027;   http:
//rmp.aps.org/abstract/RMP/v77/i3/p1027_1.

Mackenzie:2003:SSP

[MM03] Andrew Peter Mackenzie and Yoshiteru Maeno. The superconductivity of Sr_{2}RuO_{4} and the physics of spin-triplet pairing. Reviews of Modern Physics, 75(2):657–712, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.657;   http://rmp.aps.org/
abstract/RMP/v75/i2/p657_1.

Maruyama:2009:CPM

[MNV09] Koji Maruyama, Franco Nori, and Vlatko Vedral. Colloquium: The physics of Maxwell’s demon and information. Reviews of Modern Physics, 81(1):1–23, January 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.1;   http://rmp.aps.org/abstract/RMP/v81/i1/p1_1.

Morsch:2006:DBE

[MO06] Oliver Morsch and Markus Oberthaler. Dynamics of Bose– Einstein condensates in optical lattices. Reviews of Modern Physics, 78(1):179–215, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.78.179;
http://rmp.aps.org/abstract/RMP/v78/i1/p179_1.

Morgan:2007:CSB

[Mor07] Frank Morgan. Colloquium: Soap bubble clusters. Reviews of Modern Physics, 79(3):821–827, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.821;
http://rmp.aps.org/abstract/RMP/v79/i3/p821_1.

Mujica:2003:HPP

[MRMN03] A. Mujica, Angel Rubio, A. Muñoz, and R. J. Needs. High-pressure phases of group-IV, III–V, and II–VI compounds. Reviews of Modern Physics, 75(3):863–912, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.863;   http://rmp.aps.org/
abstract/RMP/v75/i3/p863_1.

Murthy:2003:HTF

[MS03] Ganpathy Murthy and R. Shankar. Hamiltonian theories of the fractional quantum Hall effect. Reviews of Modern Physics, 75 (4):1101–1158, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1101;   http:
//rmp.aps.org/abstract/RMP/v75/i4/p1101_1.

Marklund:2006:NCE

[MS06] Mattias Marklund and Padma K. Shukla. Nonlinear collective effects in photon-photon and photon-plasma interactions. Reviews of Modern Physics, 78(2):591–640, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.591;   http://rmp.aps.org/
abstract/RMP/v78/i2/p591_1.

Makhlin:2001:QSE

[MSS01] Yuriy Makhlin, Gerd Schön, and Alexander Shnirman. Quantumstate engineering with Josephson-junction devices. Reviews of Modern Physics, 73(2):357–400, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.73.357;
http://rmp.aps.org/abstract/RMP/v73/i2/p357_1.

Mohr:2000:CRV

[MT00] Peter J. Mohr and Barry N. Taylor. CODATA recommended values of the fundamental physical constants: 1998. Reviews of Modern Physics, 72(2):351–495, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.72.351;
http://rmp.aps.org/abstract/RMP/v72/i2/p351_1.

Mohr:2005:CRV

[MT05] Peter J. Mohr and Barry N. Taylor. CODATA recommended values of the fundamental physical constants: 2002. Reviews of Modern Physics, 77(1):1–107, January 2005. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
77.1;   http://rmp.aps.org/abstract/RMP/v77/i1/p1_1.

Mourou:2006:ORR

[MTB06] Gerard A. Mourou, Toshiki Tajima, and Sergei V. Bulanov. Optics in the relativistic regime. Reviews of Modern Physics, 78 (2):309–371, April 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.78.309;   http://rmp.
aps.org/abstract/RMP/v78/i2/p309_1.

Mohr:2008:CRV

[MTN08] Peter J. Mohr, Barry N. Taylor, and David B. Newell. CO-DATA recommended values of the fundamental physical constants: 2006. Reviews of Modern Physics, 80(2):633–730, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.633;   http://rmp.aps.org/
abstract/RMP/v80/i2/p633_1.

Moore:2009:NSA

[MvdL09] Kevin T. Moore and Gerrit van der Laan. Nature of the 5f states in actinide metals. Reviews of Modern Physics, 81(1): 235–298, January 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.81.235;   http://rmp.
aps.org/abstract/RMP/v81/i1/p235_1.

Maniv:2001:VSQ

[MZVW01] Tsofar Maniv, Vladimir Zhuravlev, Israel Vagner, and Peter Wyder. Vortex states and quantum magnetic oscillations in conventional type-II superconductors. Reviews of Modern Physics, 73(4):867–911, October 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.73.867;   http:
//rmp.aps.org/abstract/RMP/v73/i4/p867_1.

Nambu:2009:NLS

[Nam09] Yoichiro Nambu. Nobel Lecture: Spontaneous symmetry breaking in particle physics: a case of cross fertilization. Reviews of Modern Physics, 81(3):1015–1018, July 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.1015;   http://rmp.aps.org/abstract/RMP/v81/i3/p1015_

1. See note [Nam10].

Nambu:2010:PNN

[Nam10] Yoichiro Nambu. Publisher’s note: Nobel Lecture: Spontaneous symmetry breaking in particle physics: a case of cross fertilization [Rev. Mod. Phys. 81, 1015 (2009)]. Reviews of Modern Physics, 82(4):3199, October 2010. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.82.
3199;   http://rmp.aps.org/abstract/RMP/v82/i4/p3199_

1. See [Nam09].

Neto:2009:EPG

[NGP^{+}09] A. H. Castro Neto, F. Guinea, N. M. R. Peres, K. S. Novoselov, and A. K. Geim. The electronic properties of graphene. Reviews of Modern Physics, 81(1):109–162, January 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.81.109;   http://rmp.aps.org/
abstract/RMP/v81/i1/p109_1.

Neil:2002:TAH

[NM02] George R. Neil and Lia Merminga. Technical approaches for high-average-power free-electron lasers. Reviews of Modern Physics, 74(3):685–701, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.74.685;
http://rmp.aps.org/abstract/RMP/v74/i3/p685_1.

Nayak:2008:NAA

[NSS^{+}08] Chetan Nayak, Steven H. Simon, Ady Stern, Michael Freedman, and Sankar Das Sarma. Non-Abelian anyons and topological quantum computation. Reviews of Modern Physics, 80 (3):1083–1159, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1083;   http:
//rmp.aps.org/abstract/RMP/v80/i3/p1083_1.

Nagano:2000:OIU

[NW00] M. Nagano and A. A. Watson. Observations and implications of the ultrahigh-energy cosmic rays. Reviews of Modern Physics, 72(3):689–732, July 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.72.689;   http://rmp.
aps.org/abstract/RMP/v72/i3/p689_1.

Odor:2004:UCN

[Ódo04] Géza Ódor. Universality classes in nonequilibrium lattice systems. Reviews of Modern Physics, 76(3):663–724, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.76.663;   http://rmp.aps.org/
abstract/RMP/v76/i3/p663_1.

Ozeri:2005:CBB

[OKSD05] R. Ozeri, N. Katz, J. Steinhauer, and N. Davidson. Colloquium: Bulk Bogoliubov excitations in a Bose–Einstein condensate. Reviews of Modern Physics, 77(1):187–205, January 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.187;   http://rmp.aps.org/
abstract/RMP/v77/i1/p187_1.

Onida:2002:EED

[ORR02] Giovanni Onida, Lucia Reining, and Angel Rubio. Electronic excitations: density-functional versus many-body Green’s-function approaches. Reviews of Modern Physics, 74(2):601–659, April 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.601;   http://rmp.aps.org/
abstract/RMP/v74/i2/p601_1.

ORaifeartaigh:2000:GTH

[OS00] Lochlainn O’Raifeartaigh and Norbert Straumann. Gauge theory: Historical origins and some modern developments. Reviews of Modern Physics, 72(1):1–23, January 2000. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
72.1;   http://rmp.aps.org/abstract/RMP/v72/i1/p1_1.

Ostrikov:2005:CRP

[Ost05] K. Ostrikov. Colloquium: Reactive plasmas as a versatile nanofabrication tool. Reviews of Modern Physics, 77(2):489– 511, April 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.77.489;   http://rmp.aps.
org/abstract/RMP/v77/i2/p489_1.

Orlandini:2007:STC

[OW07] E. Orlandini and S. G. Whittington. Statistical topology of closed curves: Some applications in polymer physics. Reviews of Modern Physics, 79(2):611–642, April 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.611;
http://rmp.aps.org/abstract/RMP/v79/i2/p611_1.

Pfleiderer:2009:SPE

[Pfl09] Christian Pfleiderer. Superconducting phases of f-electron compounds. Reviews of Modern Physics, 81(4):1551–1624, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.81.1551;   http://rmp.aps.org/abstract/
RMP/v81/i4/p1551_1.

Pande:2000:HFD

[PGT00] Vijay S. Pande, Alexander Yu. Grosberg, and Toyoichi Tanaka. Heteropolymer freezing and design: Towards physical models of protein folding. Reviews of Modern Physics, 72(1):259–314, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.72.259;   http://rmp.aps.
org/abstract/RMP/v72/i1/p259_1.

Piran:2004:PGR

[Pir04] Tsvi Piran. The physics of gamma-ray bursts. Reviews of Modern Physics, 76(4):1143–1210, October 2004. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
76.1143;   http://rmp.aps.org/abstract/RMP/v76/i4/p1143_
1.

Procaccia:2008:CTD

[PLB08] Itamar Procaccia, Victor S. L’vov, and Roberto Benzi. Colloquium: Theory of drag reduction by polymers in wall-bounded turbulence. Reviews of Modern Physics, 80(1):225–247, January 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.225;   http://rmp.aps.org/
abstract/RMP/v80/i1/p225_1.

Pohl:2002:LTT

[PLT02] Robert O. Pohl, Xiao Liu, and EunJoo Thompson. Lowtemperature thermal conductivity and acoustic attenuation in amorphous solids. Reviews of Modern Physics, 74(4):991–1013, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.74.991;   http://rmp.aps.
org/abstract/RMP/v74/i4/p991_1.

Pfeifer:2007:CME

[PNHRD07] Robert N. C. Pfeifer, Timo A. Nieminen, Norman R. Heckenberg, and Halina Rubinsztein-Dunlop. Colloquium: Momentum of an electromagnetic wave in dielectric media. Reviews of Modern Physics, 79(4):1197–1216, October 2007. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
79.1197;   http://rmp.aps.org/abstract/RMP/v79/i4/p1197_

1. See erratum [PNHRD09].

Pfeifer:2009:ECM

[PNHRD09] Robert N. C. Pfeifer, Timo A. Nieminen, Norman R. Heckenberg, and Halina Rubinsztein-Dunlop. Erratum: Colloquium: Momentum of an electromagnetic wave in dielectric media [Rev. Mod. Phys. 79, 1197 (2007)]. Reviews of Modern Physics, 81 (1):443, January 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.81.443;   http://rmp.
aps.org/abstract/RMP/v81/i1/p443_1. See [PNHRD07].

Politzer:2005:NLD

[Pol05] David Politzer. Nobel Lecture: The dilemma of attribution. Reviews of Modern Physics, 77(3):851–856, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.851;   http://rmp.aps.org/
abstract/RMP/v77/i3/p851_1.

Posazhennikova:2006:CWI

[Pos06] Anna Posazhennikova. Colloquium: Weakly interacting, dilute Bose gases in 2D. Reviews of Modern Physics, 78(4): 1111–1134, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1111;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1111_1.

Peebles:2003:CCD

[PR03] P. J. E. Peebles and Bharat Ratra. The cosmological constant and dark energy. Reviews of Modern Physics, 75(2):559–606, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.75.559;   http://rmp.aps.org/
abstract/RMP/v75/i2/p559_1.

Peres:2004:QIR

[PT04] Asher Peres and Daniel R. Terno. Quantum information and Relativity theory. Reviews of Modern Physics, 76(1): 93–123, January 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.76.93;   http:/
/rmp.aps.org/abstract/RMP/v76/i1/p93_1.

Papenbrock:2007:CRM

[PW07] T. Papenbrock and H. A. Weidenmüller. Colloquium: Random matrices and chaos in nuclear spectra. Reviews of Modern Physics, 79(3):997–1013, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.997;
http://rmp.aps.org/abstract/RMP/v79/i3/p997_1.

Rehr:2000:TAX

[RA00] J. J. Rehr and R. C. Albers. Theoretical approaches to Xray absorption fine structure. Reviews of Modern Physics, 72 (3):621–654, July 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.72.621;   http://rmp.
aps.org/abstract/RMP/v72/i3/p621_1.

Raimond:2001:MQE

[RBH01] J. M. Raimond, M. Brune, and S. Haroche. Manipulating quantum entanglement with atoms and photons in a cavity. Reviews of Modern Physics, 73(3):565–582, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.565;   http://rmp.aps.org/
abstract/RMP/v73/i3/p565_1.

Reid:2009:CEP

[RDB^{+}09] M. D. Reid, P. D. Drummond, W. P. Bowen, E. G. Cavalcanti, P. K. Lam, H. A. Bachor, U. L. Andersen, and G. Leuchs. Colloquium: The Einstein–Podolsky–Rosen paradox: From concepts to applications. Reviews of Modern Physics, 81(4): 1727–1751, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1727;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1727_1.

Ryutov:2000:PFP

[RDM00] D. D. Ryutov, M. S. Derzon, and M. K. Matzen. The physics of fast Z pinches. Reviews of Modern Physics, 72(1):167–223, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.72.167;   http://rmp.aps.
org/abstract/RMP/v72/i1/p167_1.

Remington:2006:EAH

[RDR06] Bruce A. Remington, R. Paul Drake, and Dmitri D. Ryutov. Experimental astrophysics with high power lasers and Z pinches. Reviews of Modern Physics, 78(3):755–807, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.755;   http://rmp.aps.org/
abstract/RMP/v78/i3/p755_1.

Roberts:2000:GTS

[RG00] Paul H. Roberts and Gary A. Glatzmaier. Geodynamo theory and simulations. Reviews of Modern Physics, 72(4):1081– 1123, October 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.72.1081;   http:
//rmp.aps.org/abstract/RMP/v72/i4/p1081_1.

Richter:2008:HAC

[RGC^{+}08] Burton Richter, David Goldston, George Crabtree, Leon Glicksman, David Goldstein, David Greene, Dan Kammen, Mark Levine, Michael Lubell, Maxine Savitz, Daniel Sperling, Fred Schlachter, John Scofield, and James Dawson. How America can look within to achieve energy security and reduce global warming. Reviews of Modern Physics, 80(4):S1–S109, October 2008.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.80.S1;   http://rmp.aps.org/abstract/RMP/
v80/i4/pS1_1.

Rossi:2002:TUP

[RK02] Fausto Rossi and Tilmann Kuhn. Theory of ultrafast phenomena in photoexcited semiconductors. Reviews of Modern Physics, 74 (3):895–950, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.74.895;   http://rmp.
aps.org/abstract/RMP/v74/i3/p895_1.

Reimann:2002:ESQ

[RM02] Stephanie M. Reimann and Matti Manninen. Electronic structure of quantum dots. Reviews of Modern Physics, 74(4): 1283–1342, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.74.1283;   http:
//rmp.aps.org/abstract/RMP/v74/i4/p1283_1.

Rousse:2001:FXR

[RRG01] Antoine Rousse, Christian Rischel, and Jean-Claude Gauthier. Femtosecond X-ray crystallography. Reviews of Modern Physics, 73(1):17–31, January 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.73.17;   http:/
/rmp.aps.org/abstract/RMP/v73/i1/p17_1.

Rabinovich:2006:DPN

[RVSA06] Mikhail I. Rabinovich, Pablo Varona, Allen I. Selverston, and Henry D. I. Abarbanel. Dynamical principles in neuroscience. Reviews of Modern Physics, 78(4):1213– 1265, October 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.78.1213;   http:
//rmp.aps.org/abstract/RMP/v78/i4/p1213_1.

Robson:2005:CPB

[RWP05] R. E. Robson, R. D. White, and Z. Lj. Petrović. Colloquium: Physically based fluid modeling of collisionally dominated low-temperature plasmas. Reviews of Modern Physics, 77 (4):1303–1320, October 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1303;   http:
//rmp.aps.org/abstract/RMP/v77/i4/p1303_1.

Raubenheimer:2000:FFS

[RZ00] T. O. Raubenheimer and F. Zimmermann. Final-focus systems in linear colliders. Reviews of Modern Physics, 72(1): 95–107, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.72.95;   http:/
/rmp.aps.org/abstract/RMP/v72/i1/p95_1.

Sachdev:2003:COQ

[Sac03] Subir Sachdev. Colloquium: Order and quantum phase transitions in the cuprate superconductors. Reviews of Modern Physics, 75(3):913–932, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.75.913;
http://rmp.aps.org/abstract/RMP/v75/i3/p913_1.

Saha:2002:MOA

[Sah02] Swapan K. Saha. Modern optical astronomy: technology and impact of interferometry. Reviews of Modern Physics, 74(2): 551–600, April 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.74.551;   http://rmp.
aps.org/abstract/RMP/v74/i2/p551_1.

Sonier:2000:SVS

[SBK00] Jeff E. Sonier, Jess H. Brewer, and Robert F. Kiefl. μSR studies of the vortex state in type-II superconductors. Reviews of Modern Physics, 72(3):769–811, July 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.72.769;
http://rmp.aps.org/abstract/RMP/v72/i3/p769_1.

Severijns:2006:TSE

[SBNC06] Nathal Severijns, Marcus Beck, and Oscar Naviliat-Cuncic. Tests of the standard electroweak model in nuclear beta decay. Reviews of Modern Physics, 78(3):991–1040, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.78.991;   http://rmp.aps.org/
abstract/RMP/v78/i3/p991_1.

Scarani:2009:SPQ

[SBPC^{+}09] Valerio Scarani, Helle Bechmann-Pasquinucci, Nicolas J. Cerf, Miloslav Dusek, Norbert Lütkenhaus, and Momtchil Peev. The security of practical quantum key distribution. Reviews of Modern Physics, 81(3):1301–1350, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.81.
1301;   http://rmp.aps.org/abstract/RMP/v81/i3/p1301_
1.

Santini:2009:MIE

[SCA^{+}09] Paolo Santini, Stefano Carretta, Giuseppe Amoretti, Roberto Caciuffo, Nicola Magnani, and Gerard H. Lander. Multipolar interactions in f-electron systems: The paradigm of actinide dioxides. Reviews of Modern Physics, 81(2):807–863, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.81.807;   http://rmp.aps.org/
abstract/RMP/v81/i2/p807_1.

Schlosshauer:2004:DMP

[Sch04] Maximilian Schlosshauer. Decoherence, the measurement problem, and interpretations of quantum mechanics. Reviews of Modern Physics, 76(4):1267–1305, October 2004. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
76.1267;   http://rmp.aps.org/abstract/RMP/v76/i4/p1267_
1.

Schollwock:2005:DMR

[Sch05] U. Schollwöck. The density-matrix renormalization group. Reviews of Modern Physics, 77(1):259–315, January 2005. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.77.259;   http://rmp.aps.org/abstract/RMP/
v77/i1/p259_1.

Shukla:2009:CFD

[SE09] P. K. Shukla and B. Eliasson. Colloquium: Fundamentals of dust-plasma interactions. Reviews of Modern Physics, 81 (1):25–44, January 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.81.25;   http:/
/rmp.aps.org/abstract/RMP/v81/i1/p25_1.

Shankar:2008:CCQ

[Sha08] R. Shankar. Colloquium: Chaotic quantum dots with strongly correlated electrons. Reviews of Modern Physics, 80(2):379– 394, April 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.80.379;   http://rmp.aps.
org/abstract/RMP/v80/i2/p379_1.

Stangl:2004:SPS

[SHB04] J. Stangl, V. Holý, and G. Bauer. Structural properties of self-organized semiconductor nanostructures. Reviews of Modern Physics, 76(3):725–783, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.76.725;
http://rmp.aps.org/abstract/RMP/v76/i3/p725_1.

Shirakawa:2001:NLD

[Shi01] Hideki Shirakawa. Nobel Lecture: The discovery of polyacetylene film — the dawning of an era of conducting polymers. Reviews of Modern Physics, 73(3):713–718, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.713;   http://rmp.aps.org/
abstract/RMP/v73/i3/p713_1.

Schoch:2008:TPN

[SHR08] Reto B. Schoch, Jongyoon Han, and Philippe Renaud. Transport phenomena in nanofluidics. Reviews of Modern Physics, 80 (3):839–883, July 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.80.839;   http://rmp.
aps.org/abstract/RMP/v80/i3/p839_1.

Sato:2006:CNE

[SHS06] M. Sato, B. E. Hubbard, and A. J. Sievers. Colloquium: Nonlinear energy localization and its manipulation in micromechanical oscillator arrays. Reviews of Modern Physics, 78(1):137–157, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.78.137;   http://rmp.aps.
org/abstract/RMP/v78/i1/p137_1.

Sidebottom:2009:CUI

[Sid09] D. L. Sidebottom. Colloquium: Understanding ion motion in disordered solids from impedance spectroscopy scaling. Reviews of Modern Physics, 81(3):999–1014, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.81.999;   http://rmp.aps.org/
abstract/RMP/v81/i3/p999_1.

Scarani:2005:QC

[SIGA05] Valerio Scarani, Sofyan Iblisdir, Nicolas Gisin, and Antonio Acín. Quantum cloning. Reviews of Modern Physics, 77(4): 1225–1256, October 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1225;   http:
//rmp.aps.org/abstract/RMP/v77/i4/p1225_1.

Salamon:2001:PMS

[SJ01] Myron B. Salamon and Marcelo Jaime. The physics of manganites: Structure and transport. Reviews of Modern Physics, 73 (3):583–628, July 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.73.583;   http://rmp.
aps.org/abstract/RMP/v73/i3/p583_1.

Succi:2002:CRT

[SKC02] Sauro Succi, Iliya V. Karlin, and Hudong Chen. Colloquium: Role of the H theorem in lattice Boltzmann hydrodynamic simulations. Reviews of Modern Physics, 74(4): 1203–1220, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.74.1203;   http:
//rmp.aps.org/abstract/RMP/v74/i4/p1203_1.

Smoot:2007:NLC

[Smo07] George F. Smoot. Nobel Lecture: Cosmic microwave background radiation anisotropies: Their discovery and utilization. Reviews of Modern Physics, 79(4):1349–1379, October 2007.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.79.1349;   http://rmp.aps.org/abstract/RMP/
v79/i4/p1349_1.

Sprouse:2008:EAN

[Spr08a] Gene D. Sprouse. Editorial: APS now leaves copyright with authors for derivative works. Reviews of Modern Physics, 80 (4):1199, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1199;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1199_1.

Sprouse:2008:EPS

[Spr08b] Gene D. Sprouse. Editorial: Physics — spotlighting exceptional research. Reviews of Modern Physics, 80(4): 1197, October 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.80.1197;   http:
//rmp.aps.org/abstract/RMP/v80/i4/p1197_1.

Sprouse:2008:EWW

[Spr08c] Gene D. Sprouse. Editorial: Which Wei Wang? Reviews of Modern Physics, 80(1):1, January 2008. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
80.1;   http://rmp.aps.org/abstract/RMP/v80/i1/p1_1.

Squires:2005:MFP

[SQ05] Todd M. Squires and Stephen R. Quake. Microfluidics: Fluid physics at the nanoliter scale. Reviews of Modern Physics, 77 (3):977–1026, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.77.977;   http://rmp.
aps.org/abstract/RMP/v77/i3/p977_1.

Scopigno:2005:MDL

[SRS05] Tullio Scopigno, Giancarlo Ruocco, and Francesco Sette. Microscopic dynamics in liquid metals: The experimental point of view. Reviews of Modern Physics, 77(3):881–933, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.881;   http://rmp.aps.org/
abstract/RMP/v77/i3/p881_1.

Shadmi:2000:DSB

[SS00] Yael Shadmi and Yuri Shirman. Dynamical supersymmetry breaking. Reviews of Modern Physics, 72(1):25– 64, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.72.25;   http://
rmp.aps.org/abstract/RMP/v72/i1/p25_1.

Stapelfeldt:2003:CAM

[SS03] Henrik Stapelfeldt and Tamar Seideman. Colloquium: Aligning molecules with strong laser pulses. Reviews of Modern Physics, 75(2):543–557, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.75.543;   http://rmp.
aps.org/abstract/RMP/v75/i2/p543_1.

Sagues:2007:SON

[SSGO07] Francesc Sagués, José M. Sancho, and Jordi García-Ojalvo. Spatiotemporal order out of noise. Reviews of Modern Physics, 79 (3):829–882, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.79.829;   http://rmp.
aps.org/abstract/RMP/v79/i3/p829_1.

Sadri:2004:PWS

[SSJ04] Darius Sadri and M. M. Sheikh-Jabbari. The plane-wave/super Yang–Mills duality. Reviews of Modern Physics, 76(3):853– 907, July 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.76.853;   http://rmp.aps.
org/abstract/RMP/v76/i3/p853_1.

Stewart:2001:NFL

[Ste01] G. R. Stewart. Non-Fermi-liquid behavior in d- and f-electron metals. Reviews of Modern Physics, 73(4):797–855, October 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.797;   http://rmp.aps.org/
abstract/RMP/v73/i4/p797_1. See addendum [Ste06].

Stewart:2006:ANF

[Ste06] G. R. Stewart. Addendum: Non-Fermi-liquid behavior in dand f-electron metals. Reviews of Modern Physics, 78(3):743– 753, July 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.78.743;   http://rmp.aps.
org/abstract/RMP/v78/i3/p743_1. See [Ste01].

Strobl:2009:CLC

[Str09] Gert Strobl. Colloquium: Laws controlling crystallization and melting in bulk polymers. Reviews of Modern Physics, 81 (3):1287–1300, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1287;   http:
//rmp.aps.org/abstract/RMP/v81/i3/p1287_1.

Shifman:2007:SS

[SY07] M. Shifman and A. Yung. Supersymmetric solitons. Reviews of Modern Physics, 79(4):1139–1196, October 2007. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
79.1139;   http://rmp.aps.org/abstract/RMP/v79/i4/p1139_
1.

tHooft:2000:NLC

[’t 00] Gerard ’t Hooft. Nobel Lecture: a confrontation with infinity. Reviews of Modern Physics, 72(2):333–339, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.72.333;   http://rmp.aps.org/
abstract/RMP/v72/i2/p333_1. See erratum [’t 02].

tHooft:2002:ENL

[’t 02] Gerard ’t Hooft. Erratum: Nobel Lecture: A confrontation with infinity [Rev. Mod. Phys. 72, 333 (2000)]. Reviews of Modern Physics, 74(4):1343, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.74.
1343;   http://rmp.aps.org/abstract/RMP/v74/i4/p1343_

1. See [’t 00].

Taylor:2001:ATM

[Tay01] Washington Taylor. M(atrix) theory: matrix quantum mechanics as a fundamental theory. Reviews of Modern Physics, 73 (2):419–461, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.73.419;   http://rmp.
aps.org/abstract/RMP/v73/i2/p419_1.

Tserkovnyak:2005:NMD

[TBBH05] Yaroslav Tserkovnyak, Arne Brataas, Gerrit E. W. Bauer, and Bertrand I. Halperin. Nonlocal magnetization dynamics in ferromagnetic heterostructures. Reviews of Modern Physics, 77 (4):1375–1421, October 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1375;   http:
//rmp.aps.org/abstract/RMP/v77/i4/p1375_1.

Terry:2000:STT

[Ter00] P. W. Terry. Suppression of turbulence and transport by sheared flow. Reviews of Modern Physics, 72(1):109–165, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.72.109;   http://rmp.aps.org/
abstract/RMP/v72/i1/p109_1.

Teubner:2009:HOH

[TG09] U. Teubner and P. Gibbon. High-order harmonics from laserirradiated plasma surfaces. Reviews of Modern Physics, 81(2): 445–479, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.81.445;   http://rmp.
aps.org/abstract/RMP/v81/i2/p445_1.

Thoma:2009:CFT

[Tho09] Markus H. Thoma. Colloquium: Field theoretic description of ultrarelativistic electron-positron plasmas. Reviews of Modern Physics, 81(3):959–968, July 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.959;
http://rmp.aps.org/abstract/RMP/v81/i3/p959_1.

Tsuei:2000:PSC

[TK00] C. C. Tsuei and J. R. Kirtley. Pairing symmetry in cuprate superconductors. Reviews of Modern Physics, 72(4):969–1016, October 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.72.969;   http://rmp.aps.
org/abstract/RMP/v72/i4/p969_1.

Thurman-Keup:2001:WBP

[TKKTBW01] Randy M. Thurman-Keup, Ashutosh V. Kotwal, Monica Tecchio, and Aesook Byon-Wagner. W boson physics at hadron colliders. Reviews of Modern Physics, 73(2):267–306, April 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.73.267;   http://rmp.aps.org/
abstract/RMP/v73/i2/p267_1.

Tanner:2000:TTE

[TRR00] Gregor Tanner, Klaus Richter, and Jan-Michael Rost. The theory of two-electron atoms: between ground state and complete fragmentation. Reviews of Modern Physics, 72(2):497– 544, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.72.497;   http://rmp.aps.
org/abstract/RMP/v72/i2/p497_1.

Tsori:2009:CPT

[Tso09] Yoav Tsori. Colloquium: Phase transitions in polymers and liquids in electric fields. Reviews of Modern Physics, 81(4): 1471–1494, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1471;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1471_1.

Uggerhoj:2005:IRP

[Ugg05] Ulrik I. Uggerhøj. The interaction of relativistic particles with strong crystalline fields. Reviews of Modern Physics, 77(4): 1131–1171, October 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.77.1131;   http:
//rmp.aps.org/abstract/RMP/v77/i4/p1131_1.

Uzan:2003:FCT

[Uza03] Jean-Philippe Uzan. The fundamental constants and their variation: observational and theoretical status. Reviews of Modern Physics, 75(2):403–455, April 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.75.403;
http://rmp.aps.org/abstract/RMP/v75/i2/p403_1.

Vandersypen:2004:NTQ

[VC04] L. M. K. Vandersypen and I. L. Chuang. NMR techniques for quantum control and computation. Reviews of Modern Physics, 76(4):1037–1069, October 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.76.1037;   http:
//rmp.aps.org/abstract/RMP/v76/i4/p1037_1.

vandeWalle:2002:ELV

[vdWC02] A. van de Walle and G. Ceder. The effect of lattice vibrations on substitutional alloy thermodynamics. Reviews of Modern Physics, 74(1):11–45, January 2002. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
74.11;   http://rmp.aps.org/abstract/RMP/v74/i1/p11_
1.

vanderWiel:2003:ETT

[vdWDE^{+}03] W. G. van der Wiel, S. De Franceschi, J. M. Elzerman, T. Fujisawa, S. Tarucha, and L. P. Kouwenhoven. Electron transport through double quantum dots. Reviews of Modern Physics, 75(1):1–22, January 2003. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
75.1;   http://rmp.aps.org/abstract/RMP/v75/i1/p1_1.

Vedral:2002:RRE

[Ved02] V. Vedral. The role of relative entropy in quantum information theory. Reviews of Modern Physics, 74(1):197–234, January 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.197;   http://rmp.aps.org/
abstract/RMP/v74/i1/p197_1.

Veltman:2000:NLW

[Vel00] Martinus J. G. Veltman. Nobel Lecture: From weak interactions to gravitation. Reviews of Modern Physics, 72(2):341– 349, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.72.341;   http://rmp.aps.
org/abstract/RMP/v72/i2/p341_1.

Veysey:2007:SVF

[VG07] John Veysey II and Nigel Goldenfeld. Simple viscous flows: From boundary layers to the renormalization group. Reviews of Modern Physics, 79(3):883–927, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.883;
http://rmp.aps.org/abstract/RMP/v79/i3/p883_1.

Viovy:2000:EDO

[Vio00] Jean-Louis Viovy. Electrophoresis of DNA and other polyelectrolytes: Physical mechanisms. Reviews of Modern Physics, 72 (3):813–872, July 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.72.813;   http://rmp.
aps.org/abstract/RMP/v72/i3/p813_1.

Lohneysen:2007:FLI

[vLRVW07] Hilbert v. Löhneysen, Achim Rosch, Matthias Vojta, and Peter Wölfle. Fermi-liquid instabilities at magnetic quantum phase transitions. Reviews of Modern Physics, 79(3): 1015–1075, July 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.79.1015;   http:
//rmp.aps.org/abstract/RMP/v79/i3/p1015_1.

Voit:2005:TCE

[Voi05] G. Mark Voit. Tracing cosmic evolution with clusters of galaxies. Reviews of Modern Physics, 77(1):207–258, January 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.77.207;   http://rmp.aps.org/
abstract/RMP/v77/i1/p207_1.

Volokitin:2007:NFR

[VP07] A. I. Volokitin and B. N. J. Persson. Near-field radiative heat transfer and noncontact friction. Reviews of Modern Physics, 79 (4):1291–1329, October 2007. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.79.1291;   http:
//rmp.aps.org/abstract/RMP/v79/i4/p1291_1.

vanTiggelen:2000:NLC

[vTS00] Bart van Tiggelen and Holger Stark. Nematic liquid crystals as a new challenge for radiative transfer. Reviews of Modern Physics, 72(4):1017–1039, October 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.72.1017;   http:
//rmp.aps.org/abstract/RMP/v72/i4/p1017_1.

Wei:2003:SAH

[Wei03] Jie Wei. Synchrotrons and accumulators for high-intensity proton beams. Reviews of Modern Physics, 75(4):1383– 1432, October 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.75.1383;   http:
//rmp.aps.org/abstract/RMP/v75/i4/p1383_1.

Willmott:2000:PLV

[WH00] P. R. Willmott and J. R. Huber. Pulsed laser vaporization and deposition. Reviews of Modern Physics, 72(1):315–328, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.72.315;   http://rmp.aps.
org/abstract/RMP/v72/i1/p315_1.

Weichman:2001:CSL

[WHG01] Peter B. Weichman, Alexa W. Harter, and David L. Goodstein. Criticality and superfluidity in liquid ^{4}He under nonequilibrium conditions. Reviews of Modern Physics, 73(1):1–15, January 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.73.1;   http://rmp.aps.org/abstract/
RMP/v73/i1/p1_1.

Woosley:2002:EEM

[WHW02] S. E. Woosley, A. Heger, and T. A. Weaver. The evolution and explosion of massive stars. Reviews of Modern Physics, 74 (4):1015–1071, October 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.74.1015;   http:
//rmp.aps.org/abstract/RMP/v74/i4/p1015_1.

Widrow:2002:OGE

[Wid02] Lawrence M. Widrow. Origin of galactic and extragalactic magnetic fields. Reviews of Modern Physics, 74(3):775–823, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.775;   http://rmp.aps.org/
abstract/RMP/v74/i3/p775_1.

Wiesendanger:2009:SMN

[Wie09] Roland Wiesendanger. Spin mapping at the nanoscale and atomic scale. Reviews of Modern Physics, 81(4):1495– 1550, October 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.81.1495;   http:
//rmp.aps.org/abstract/RMP/v81/i4/p1495_1.

Willenbrock:2000:STQ

[Wil00] S. Willenbrock. Studying the top quark. Reviews of Modern Physics, 72(4):1141–1148, October 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.72.
1141;   http://rmp.aps.org/abstract/RMP/v72/i4/p1141_
1.

Wilczek:2005:NLA

[Wil05] Frank Wilczek. Nobel Lecture: Asymptotic freedom: From paradox to paradigm. Reviews of Modern Physics, 77(3):857– 870, July 2005. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.77.857;   http://rmp.aps.
org/abstract/RMP/v77/i3/p857_1.

Witten:2007:SFE

[Wit07] T. A. Witten. Stress focusing in elastic sheets. Reviews of Modern Physics, 79(2):643–675, April 2007. CODEN RMPHAT.

ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.79.643;
http://rmp.aps.org/abstract/RMP/v79/i2/p643_1.

Weidenmuller:2009:RMC

[WM09] H. A. Weidenmüller and G. E. Mitchell. Random matrices and chaos in nuclear physics: Nuclear structure. Reviews of Modern Physics, 81(2):539–589, April 2009. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.81.539;
http://rmp.aps.org/abstract/RMP/v81/i2/p539_1.

Winterfeldt:2008:COC

[WSG08] Carsten Winterfeldt, Christian Spielmann, and Gustav Gerber. Colloquium: Optimal control of high-harmonic generation. Reviews of Modern Physics, 80(1):117–140, January 2008. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.80.117;   http://rmp.aps.org/
abstract/RMP/v80/i1/p117_1.

Weisse:2006:KPM

[WWAF06] Alexander Weiße, Gerhard Wellein, Andreas Alvermann, and Holger Fehske. The kernel polynomial method. Reviews of Modern Physics, 78(1):275–306, January 2006. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.78.275;
http://rmp.aps.org/abstract/RMP/v78/i1/p275_1.

Yakovenko:2009:CSM

[YR09] Victor M. Yakovenko and J. Barkley Rosser, Jr. Colloquium: Statistical mechanics of money, wealth, and income. Reviews of Modern Physics, 81(4):1703–1725, October 2009. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
81.1703;   http://rmp.aps.org/abstract/RMP/v81/i4/p1703_
1.

Zandvliet:2000:ES

[Zan00] H. J. W. Zandvliet. Energetics of Si(001). Reviews of Modern Physics, 72(2):593–602, April 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL

http://link.aps.org/doi/10.1103/RevModPhys.72.593;

http://rmp.aps.org/abstract/RMP/v72/i2/p593_1. See er-

ratum [Zan01].

Zandvliet:2001:EES

[Zan01] H. J. W. Zandvliet. Erratum: Energetics of Si(001) [Rev. Mod. Phys. 72, 593 (2000)]. Reviews of Modern Physics, 73(1):247– 462, January 2001. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.73.247;   http://rmp.aps.
org/abstract/RMP/v73/i1/p247_1. See [Zan00].

Zwolak:2008:CPA

[ZD08] Michael Zwolak and Massimiliano Di Ventra. Colloquium: Physical approaches to DNA sequencing and detection. Reviews of Modern Physics, 80(1):141–165, January 2008. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
80.141;   http://rmp.aps.org/abstract/RMP/v80/i1/p141_
1.

Zdesenko:2002:CFD

[Zde02] Yuri Zdesenko. Colloquium: The future of double β decay research. Reviews of Modern Physics, 74(3):663–684, July 2002. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/

doi/10.1103/RevModPhys.74.663;   http://rmp.aps.org/
abstract/RMP/v74/i3/p663_1.

Zutic:2004:SFA

[ZFS04] Igor Zutić, Jaroslav Fabian, and S. Das Sarma. Spintronics: Fundamentals and applications. Reviews of Modern Physics, 76 (2):323–410, April 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.76.323;   http://rmp.
aps.org/abstract/RMP/v76/i2/p323_1.

Zhou:2004:CMT

[ZMD04] Ye Zhou, W. H. Matthaeus, and P. Dmitruk. Colloquium: Magnetohydrodynamic turbulence and time scales in astrophysical and space plasmas. Reviews of Modern Physics, 76(4): 1015–1035, October 2004. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.76.1015;   http:
//rmp.aps.org/abstract/RMP/v76/i4/p1015_1.

Zurek:2003:DEQ

[Zur03] Wojciech Hubert Zurek. Decoherence, einselection, and the quantum origins of the classical. Reviews of Modern Physics, 75(3):715–775, July 2003. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.75.715;   http://rmp.
aps.org/abstract/RMP/v75/i3/p715_1.